### v2

# Divide, Interact, Sample: The Two-System Paradigm

Abstract
Mean-field, ensemble-chain, and adaptive samplers have historically been viewed as distinct approaches to Monte Carlo sampling. In this paper, we present a unifying two-system framework that brings all three under one roof. In our approach, an ensemble of particles is split into two interacting subsystems that propose updates for each other in a symmetric, alternating fashion. This cross-system interaction ensures that the overall ensemble has $\rho(x)$ as its invariant distribution in both the finite-particle setting and the mean-field limit. The two-system construction reveals that ensemble-chain samplers can be interpreted as finite-$N$ approximations of an ideal mean-field sampler; conversely, it provides a principled recipe to discretize mean-field Langevin dynamics into tractable parallel MCMC algorithms. The framework also connects naturally to adaptive single-chain methods: by replacing particle-based statistics with time-averaged statistics from a single chain, one recovers analogous adaptive dynamics in the long-time limit without requiring a large ensemble. We derive novel two-system versions of both overdamped and underdamped Langevin MCMC samplers within this paradigm. Across synthetic benchmarks and real-world posterior inference tasks, these two-system samplers exhibit significant performance gains over the popular No-U-Turn Sampler, achieving an order of magnitude higher effective sample sizes per gradient evaluation.
Keywords Adaptive Samplers $\cdot$ Ensemble Chain Samplers $\cdot$ Langevin $\cdot$ Mean-Field Samplers
1 Introduction
Sampling from high-dimensional probability distributions is pivotal in modern Bayesian statistics [35, 44, 45]. Analytical formulas for moments or normalizing constants are often infeasible to derive or compute directly in such settings. Monte Carlo sampling provides a practical alternative: by drawing representative samples, one can approximate the expectation of a posterior while incorporating prior information, and also quantify uncertainty via posterior intervals [12, 17].
Two broad classes of MCMC algorithms are widely used: (i) Discrete-time methods based on the Metropolis--Hastings algorithm, and (ii) Continuous-time methods defined by Langevin-type stochastic differential equations (SDEs). Both approaches yield Markov chains with the desired target distribution as their stationary distribution. The increasing availability of parallel computing has made it practical to run multiple chains simultaneously, which in turn has spurred the development of ensemble-based methods that leverage interactions among chains to improve convergence.
Modern sampling strategies often adapt their proposal mechanism using either information from other chains or from the chain’s own history: Ensemble-chain samplers update each chain using statistics computed from a subset of the other chains. Mean-field samplers extend this idea to the infinite-particle regime, letting each particle use information from the entire ensemble. These methods only converge to the target distribution as the number of particles approaches infinity. Adaptive samplers use a single chain and adjust its proposal parameters on the fly based on the chain’s past samples (e.g., continually tuning the proposal covariance based on a running estimate). This adaptation breaks the Markov property, but convergence to the target can still be established under suitable conditions.
To unify these perspectives, we introduce the two-system approach. This framework divides an ensemble of $N$ particles into two interacting subsystems, with each subsystem using the empirical distribution of the other as the basis for its proposals. This symmetric coupling guarantees that the target distribution $\rho(x)$ is invariant in both the finite-$N$ setting and the infinite-particle (mean-field) limit. Starting from a mean-field Langevin SDE, we show that the two-system formulation naturally yields ensemble-chain samplers via finite-particle discretization of the continuous dynamics. By alternating updates between the two subsystems in discrete time (with appropriate Metropolis--Hastings corrections), we obtain parallel MCMC algorithms that preserve detailed balance with respect to $\rho$ and are well-suited for modern multi-core hardware.
This two-system perspective also clarifies the relationship between ensemble and mean-field methods. At each iteration, an ensemble-chain sampler uses finite-sample estimates (e.g., a sample covariance) in place of exact population quantities; as $N\to\infty$, these estimates converge to their mean-field counterparts and the particle distribution approaches $\rho$. In the long-time limit, the particles approximate independent draws from $\rho$, and the empirical statistics converge to the corresponding expectations under $\rho$. Similarly, an adaptive single-chain sampler can be viewed as the $N=1$ analog of a mean-field scheme: instead of averaging over many particles at a given time, it averages over many iterations of one chain. As the number of iterations grows, the running time-average of the statistic converges to the true expectation under $\rho$. In this way, adaptive samplers mirror the effect of ensemble methods, substituting temporal averaging for population averaging.
In summary, this work makes the following contributions:
  • We propose a unified two-system sampling framework that encompasses ensemble-chain MCMC, mean-field samplers, and adaptive samplers within a single coherent paradigm.
  • Using this framework, we derive novel two-system MCMC algorithms for both overdamped and underdamped Langevin dynamics. Specifically, two-system versions of the Metropolis-Adjusted Langevin Algorithm (MALA) and the Metropolis-Adjusted Kinetic Langevin Algorithm (MAKLA), in both ensemble-chain and adaptive forms.
  • Through extensive experiments, we provide empirical evidence of significant performance gains over state-of-the-art methods. In particular, our adaptive two-system samplers outperform their non-adaptive counterparts by up to two orders of magnitude on challenging synthetic targets, and they achieve roughly an order-of-magnitude higher efficiency (in median ESS per gradient evaluation) than the No-U-Turn Sampler (NUTS) on a diverse collection of real-world posterior inference problems.
Notation: Throughout the paper, $\rho:\mathbb{R}^{d}\to[0,\infty)$ denotes the target probability density we wish to sample from. We assume $\rho(x)$ is bounded, absolutely continuous, and that $\log\rho(x)$ is $C^{2}$-smooth. We let $S^{d}_{++}$ denote the set of real $d\times d$ symmetric positive-definite matrices, $I_{d}$ the $d\times d$ identity matrix, $\mathcal{P}(\mathbb{R}^{d})$ the space of probability distributions on $\mathbb{R}^{d}$ (with densities with respect to Lebesgue measure), and $\mathcal{B}(\mathbb{R}^{d})$ the Borel $\sigma$-algebra on $\mathbb{R}^{d}$.
2 Related Work and Paper Overview
Our work lies at the intersection of ensemble-chain MCMC, adaptive sampling, and mean-field dynamics. Prior methods have explored various combinations of these ideas, but we provide a systematic two-system framework that unifies and extends these lines of work with minimal computational overhead.
The well-known affine invariant ensemble sampler was proposed by Goodman and Weare [13]. This is an efficient sampler low-dimensional problems, but it does not scale as well in higher dimensions as Langevin-based methods.
Overdamped Langevin Samplers. Garbuno-Inigo et al. [11] proposed a mean-field ensemble overdamped Langevin SDE that converges to the correct target only in the infinite-particle (mean-field) limit. Nüsken and Reich [32] later corrected the finite-$N$ invariant measure using a leave-one-out scheme, but at the cost of $\mathcal{O}(N)$ covariance computations per step. In contrast, our two-system formulation (Section 4) achieves the same correction using only two covariance evaluations per step (independent of $N$). We further present finite-particle discretizations (Section 5.1) that preserve the target distribution without requiring the mean-field limit.
Underdamped Langevin Samplers. Recent works [25, 36, 15] have introduced ensemble-based Metropolis-adjusted underdamped Langevin samplers that adapt the momentum-refreshment step, following strategies similar to ensemble Hamiltonian Monte Carlo (HMC) [6]. However, these adaptations are limited to heuristic changes in the discretization and do not modify the underlying SDE itself. By contrast, our approach directly adapts the continuous-time dynamics via a preconditioned underdamped Langevin process in the two-system framework. Additionally, whereas the above methods adjust the integration step size to cope with ill-conditioning, we mitigate poor conditioning through a randomized step size (Section 6.2) and a one-time preprocessing step (Section 6.4) that rescales the target based on local curvature.
Paper Overview. Section 3 reviews necessary background on Langevin dynamics, mean-field MCMC, and adaptive samplers. Section 4 introduces the two-system framework at the level of interacting particle systems and their mean-field limits. Section 5 describes how this framework yields a general recipe for constructing two-system samplers, including connections to adaptive and ensemble-based updates. Section 6 develops concrete Metropolis-adjusted algorithms for overdamped and underdamped Langevin dynamics. Section 7 presents empirical results across challenging targets, with comparisons to ensemble and adaptive baselines.
3 Background
As mentioned above, samplers can broadly be categorized into continuous-time methods (based on Langevin-type SDEs) and discrete-time methods (e.g., the Metropolis--Hastings algorithm). Continuous-time samplers leverage the fact that a Langevin SDE has the target density $\rho(x)$ as its stationary distribution, meaning that as $t\to\infty$, the law of the SDE converges to $\rho(x)$. However, in practice, one must discretize the SDE to simulate it, and such discretizations typically introduce bias and no longer preserve $\rho(x)$ exactly. To correct this bias, the discretized SDE can be used as a proposal within a Metropolis--Hastings framework, yielding a Markov chain that converges to the correct target. More broadly, the Metropolis--Hastings algorithm provides a general recipe for constructing discrete-time samplers that converge to $\rho(x)$. In high-dimensional settings, however, naive proposals can lead to very low acceptance rates, resulting in poor mixing and inefficient exploration of the target.
3.1 Discrete-Time Samplers
The classic Metropolis--Hastings algorithm [14, 30] constructs a Markov chain $\{X_{k}\}_{k\geq 0}$ with stationary distribution $\rho(x)$. Each iteration consists of two steps: (i) Given the current state $X_{k}$, propose a new state $Y_{k}\sim Q_{C}(\cdot\mid X_{k})$, where $Q_{C}(\cdot\mid X_{k})$ is a proposal distribution parameterized by some matrix $C$ (e.g., a covariance); (ii) Accept the proposal with probability
$A(X_{k},Y_{k})\ =\ \min\!\Big(1,\ \frac{\rho(Y_{k})}{\rho(X_{k})}\,\frac{Q_{C}(X_{k}\mid Y_{k})}{Q_{C}(Y_{k}\mid X_{k})}\Big),$ (1)
and set $X_{k+1}=Y_{k}$ if accepted (otherwise $X_{k+1}=X_{k}$). We refer to [38, 39, 40] for more details on conditions ensuring detailed balance (and hence convergence) for the proposal distribution $Q_{C}$ and acceptance ratio $A$.
Mean-field samplers [10] modify the proposal distribution to depend on a summary statistic of the current law of the chain. Let $\theta:\mathbb{R}^{d}\to\mathbb{R}^{m}$ be a bounded, measurable, continuous function. The proposal at iteration $k$ is written as
$Q_{\Theta(\mu_{k})}(\cdot\mid X_{k}),\qquad\text{where}\qquad\Theta(\mu_{k})\ =\ \int\theta(x)\,\mu_{k}(dx),\quad\text{and}\quad\mu_{k}=\textrm{Law}(X_{k}).$ (2)
For example, if $\Theta(\mu_{k})$ is the covariance matrix of $\mu_{k}$, one might define $Q_{\Theta(\mu_{k})}(y\mid X_{k})$ to be a Gaussian proposal centered at $X_{k}$ with covariance $\Theta(\mu_{k})$.
Since $\mu_{k}$ is generally unknown in closed form, one can approximate it using an empirical measure from $N$ particles. Suppose we simulate $N$ particles $X_{k}^{1},\ldots,X_{k}^{N}$ (approximately i.i.d. from $\mu_{k}$), and define the empirical distribution
$\delta_{\mathbf{X}_{k}}\ =\ \frac{1}{N}\sum_{i=1}^{N}\delta_{X_{k}^{i}},$ (3)
where $\delta_{x}$ denotes a Dirac at $x$. The finite-particle approximation then evolves the particles via
$X_{k+1}^{i}\ \sim\ P_{\delta_{\mathbf{X}_{k}}}(X_{k}^{i}\to Y_{k}^{i}),\qquad\text{with}\qquad Y_{k}^{i}\sim Q_{\Theta(\delta_{\mathbf{X}_{k}})}(\cdot\mid X_{k}^{i}),\quad i\in\{1,\ldots,N\}.$ (4)
The infinite-particle limit of this scheme recovers the mean-field dynamics (2), for which $\rho$ is the invariant measure. However, for any fixed finite $N$, the particle system generally does not have $\rho$ as its stationary distribution [41].
Ensemble-chain methods evolve multiple interacting Markov chains simultaneously. Let $\mathbf{X}_{k}=\{X_{k}^{i}\}_{i=1}^{N}$ denote the collection of $N$ chains at time $k$. We define a data-dependent statistic based on a designated subset $S\subset\{1,\ldots,N\}$ of “reference” particles:
$\Theta(S)\ =\ \frac{1}{|S|}\sum_{i\in S}\theta(X_{k}^{i})\,$, (5)
where $S$ indexes the subset of particles that will be held fixed during the next update. (Equivalently, we can view $\Theta$ as a functional that can act either on a distribution or on a set of sample points.) Using this statistic, we propose updates for the complementary set $\{1,\ldots,N\}\setminus S$ as
$X_{k+1}^{i}\sim P_{\Theta(S)}(X_{k}^{i}\to Y_{k}^{i}),\qquad\text{with}\qquad Y_{k}^{i}\sim Q_{\Theta(S)}(\cdot\mid X_{k}^{i}),\quad\forall\,i\in\{1,\ldots,N\}\setminus S,$ (6)
and we set $X_{k+1}^{i}=X_{k}^{i}$ for all $i\in S$. Importantly, keeping the particles in $S$ fixed (i.e., excluding self-interaction) is essential to ensure that the joint state $\mathbf{X}_{k}$ has $\rho^{\otimes N}$ as its invariant distribution [32]. This exclusion principle mirrors the leave-one-out idea in continuous-time ensemble samplers (see Section 3.2) and guarantees detailed balance for the combined ensemble update.
Adaptive samplers use a single chain but update proposal parameters “on the fly” based on the chain’s history. For example, one can maintain a running estimate of $\Theta$ recursively as
$\Theta_{K}\ =\ \frac{1}{K}\theta(X_{K})\ +\ \Big(1-\frac{1}{K}\Big)\Theta_{K-1}\ =\ \frac{1}{K}\sum_{k=1}^{K}\theta(X_{k})\,$, (7)
and then take the proposal distribution to be $Q_{\Theta_{K}}(\cdot\mid X_{K})$. Because the proposal now depends on past samples, the resulting process is no longer Markovian. To ensure convergence to $\rho$, two key conditions are typically required:
  • Geometric ergodicity. For every fixed parameter value $\gamma\in\Gamma$,
    $\lVert P_{\gamma}^{n}(x,\cdot)-\rho(\cdot)\rVert_{\text{TV}}\ \leq\ c\,r^{n},$
    for some constants $c>0$, $0<r<1$. Here $P^{n}_{\gamma}$ is the $n$-step transition kernel and $\|\cdot\|_{\text{TV}}$ denotes the total variation norm. Intuitively, this says that if the algorithm is run without adaptation (parameter held fixed at $\gamma$), it should converge geometrically fast to the target.
  • Diminishing adaptation.
    $\lim_{n\to\infty}\ \sup_{x\in\mathbb{R}^{d}}\,\lVert P_{\gamma_{n+1}}(x,\cdot)-P_{\gamma_{n}}(x,\cdot)\rVert_{\text{TV}}\ =\ 0,$
    in probability. This guarantees that the magnitude of adaptation vanishes as the chain progresses.
Under these conditions (plus some additional technical assumptions), the adaptive algorithm remains ergodic and satisfies a law of large numbers: for any bounded measurable test function $h:\mathbb{R}^{d}\to\mathbb{R}$,
$\frac{1}{K}\sum_{k=1}^{K}h(X_{k})\ \to\ \int h(x)\,\rho(x)\,dx,$ (8)
in probability as $K\to\infty$ [27]. (A strong law of large numbers can also be established under stronger assumptions on $\rho$ and the kernels [22].) For more details on adaptive MCMC, we refer the reader to [27, 22].
3.2 Continuous-Time Samplers
Continuous-time samplers evolve a stochastic process whose stationary law is the target distribution $\rho$. The most common examples are based on the overdamped and underdamped Langevin dynamics. The overdamped Langevin SDE [23, 26] is given by
$dX_{t}=C_{0}\,\nabla\log\rho(X_{t})\,dt\;+\;\sqrt{2\,C_{0}}\,dW_{t}\,$, (9)
where $C_{0}\in S^{d}_{++}$ (often $C_{0}=I_{d}$), $X_{t}\in\mathbb{R}^{d}$, and $W_{t}$ is standard $d$-dimensional Brownian motion. The underdamped Langevin (kinetic Langevin) dynamics introduces an auxiliary velocity variable $V_{t}$ and evolves as
[…]
with constants $\alpha,\gamma>0$ and initial conditions $X_{0},V_{0}\in\mathbb{R}^{d}$.
Under our assumptions on $\rho(x)$, if $\rho$ further satisfies a strong log-concavity condition (i.e., $\log\rho(x)$ has a globally bounded Hessian), then the law of $X_{t}$ under (9) converges to $\rho$, and the law of $(X_{t},V_{t})$ under (10) converges to the product measure $\rho(x)\,g(v)$, where $g(v)$ is the standard Gaussian density on velocities [34].
Mean-Field Sampler (Continuous-Time). Garbuno-Inigo et al. [11] proposed a mean-field variant of the overdamped Langevin SDE. This sampler evolves the nonlinear SDE
$dX_{t}=C(\mu_{t})\,\nabla\log\rho(X_{t})\,dt\;+\;\sqrt{2\,C(\mu_{t})}\,dW_{t},$ (11)
where $\mu_{t}=\textrm{Law}(X_{t})$ and $C(\mu_{t})$ denotes (for instance) the covariance matrix of $\mu_{t}$. An analogous mean-field extension exists for the underdamped Langevin case. Under appropriate conditions, the law of $X_{t}$ in (11) still converges to $\rho$ as $t\to\infty$.
As in the discrete-time setting, to simulate (11) one must use a finite-particle approximation, evolving particles according to
$dX_{t}^{i}=C(\delta_{\mathbf{X}_{t}})\,\nabla\log\rho(X_{t}^{i})\,dt\;+\;\sqrt{2\,C(\delta_{\mathbf{X}_{t}})}\,dW_{t}^{i},\qquad i\in\{1,\ldots,N\},$ (12)
with each $W_{t}^{i}$ an independent Brownian motion.
Ensemble-Chain Sampler (Continuous-Time). Nüsken and Reich [32] observed that the naive particle approximation above does not preserve $\rho^{\otimes N}$ as an invariant joint law. They proposed the modified dynamics
$dX_{t}^{i}\ =\ C(\{1,\ldots,N\}\setminus\{i\})\,\nabla\log\rho(X_{t}^{i})\,dt\;+\;\sqrt{2\,C(\{1,\ldots,N\}\setminus\{i\})}\,dW_{t}^{i},\qquad i\in\{1,\ldots,N\}\,$, (13)
where $C(\{1,\ldots,N\}\setminus\{i\})$ denotes the covariance computed over all particles except the $i$th. This exclusion of self-interaction guarantees that $\rho^{\otimes N}$ remains invariant, directly paralleling the discrete-time case.
Adaptive Samplers (Continuous-Time). Although less explored, adaptive Langevin dynamics have been proposed in recent work [19, 24]. These methods dynamically modify aspects of the SDE (typically the step size or preconditioner) based on the trajectory’s history. In this paper, we focus on discrete-time adaptive samplers, which are more straightforward to analyze theoretically and implement in practice.
4 A Two-System Approach for McKean--Vlasov Equations
We now develop the two-system approach for general continuous-time mean-field SDEs (McKean--Vlasov equations), of which the overdamped and underdamped Langevin equations are special cases. The key idea is to consider two coupled systems of particles such that each system “drives” the evolution of the other. In a Langevin setting, this symmetric coupling is constructed to preserve the target distribution and to directly connect the behavior of the finite-$N$ particle system with its $N\to\infty$ mean-field limit.
The mean-field Langevin equation (11) is an example of a more general McKean--Vlasov SDE, which takes the form:
$d\overline{X}_{t}\ =\ b(t,\overline{X}_{t},\mu_{t})\,dt\ +\ \sigma(t,\overline{X}_{t},\mu_{t})\,dW_{t},\qquad t\in[0,T],$ (14)
where $\mu_{t}=\textrm{Law}(\overline{X}_{t})$, and $b:[0,T]\times\mathbb{R}^{d}\times\mathcal{P}(\mathbb{R}^{d})\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\times\mathcal{P}(\mathbb{R}^{d})\to\mathbb{R}^{d\times d}$ are given drift and diffusion functions. To ensure existence and uniqueness of strong solutions [7, 42], we assume $b$ and $\sigma$ satisfy:
Assumption 1 (Finite second moments).
For finite $T$,
$\mathbb{E}\int_{0}^{T}\lvert b(t,x,\mu)\rvert^{2}\,dt<\infty,\qquad\mathbb{E}\int_{0}^{T}\lvert\sigma(t,x,\mu)\rvert^{2}\,dt<\infty,$ (15)
for the relevant processes (with norms understood in Euclidean and Frobenius sense, respectively).
Assumption 2 (Lipschitz continuity).
There exists a constant $C>0$ such that for all $t\in[0,T]$, $x_{1},x_{2}\in\mathbb{R}^{d}$, and $\mu_{1},\mu_{2}\in\mathcal{P}(\mathbb{R}^{d})$,
$\lvert b(t,x_{1},\mu_{1})-b(t,x_{2},\mu_{2})\rvert\ +\ \lvert\sigma(t,x_{1},\mu_{1})-\sigma(t,x_{2},\mu_{2})\rvert$ (16)
$\qquad\qquad\leq\ C\,\big(\lvert x_{1}-x_{2}\rvert+\mathcal{W}_{2}(\mu_{1},\mu_{2})\big)\,$,
where $\mathcal{W}_{2}$ denotes the 2-Wasserstein distance.
The classical particle approximation to (14) is given by the system of $N$ interacting SDEs:
$dX_{t}^{j}\ =\ b(t,X_{t}^{j},\delta_{\mathbf{X}_{t}})\,dt\ +\ \sigma(t,X_{t}^{j},\delta_{\mathbf{X}_{t}})\,dW_{t}^{j},\qquad j=1,\dots,N,$ (17)
with $W_{t}^{1},\ldots,W_{t}^{N}$ independent Brownian motions and $\mathbf{X}_{t}=\{X_{t}^{j}\}_{j=1}^{N}$. As $N\to\infty$, this interacting particle system converges to the mean-field equation (14) in the sense of propagation of chaos [9, 42]: the particles become asymptotically independent and each follows the law of the McKean--Vlasov solution $\overline{X}_{t}$.
The Two-System Approach.
Following the two-system paradigm, we introduce a coupled SDE on an augmented state $Z_{t}=(Z_{1,t},Z_{2,t})\in\mathbb{R}^{2d}$, where $Z_{1,t},Z_{2,t}\in\mathbb{R}^{d}$. We define
$dZ_{t}\ =\ \mathbf{b}(t,Z_{t},\pi_{t})\,dt\ +\ \bm{\sigma}(t,Z_{t},\pi_{t})\,dW_{t},$ (18)
where $\pi_{t}=\textrm{Law}(Z_{t})$ and $W_{t}$ is a Brownian motion in $\mathbb{R}^{2d}$. The drift and diffusion are constructed as
$\mathbf{b}(t,Z,\pi)\ =\ \begin{pmatrix}b(t,Z_{1},\,p_{2}\!\circ\pi)\\[6.0pt] b(t,Z_{2},\,p_{1}\!\circ\pi)\end{pmatrix},\qquad\bm{\sigma}(t,Z,\pi)\ =\ \begin{pmatrix}\sigma(t,Z_{1},\,p_{2}\!\circ\pi)&0\\[6.0pt] 0&\sigma(t,Z_{2},\,p_{1}\!\circ\pi)\end{pmatrix}$, (19)
with $p_{1}$ and $p_{2}$ denoting the projection maps from a distribution on $\mathbb{R}^{2d}$ to its first and second $\mathbb{R}^{d}$-marginals, respectively. In words, the two-system SDE (18) consists of two processes $Z_{1,t}$ and $Z_{2,t}$, each evolving according to the original drift/diffusion coefficients $b,\sigma$ evaluated at the empirical law of the other system.
Under Assumptions 1 and 2, one can show that the coupled SDE (18) has a unique strong solution on $[0,T]$ (proof in Appendix A.2). Moreover, $\textrm{Law}(Z_{1,t})$ and $\textrm{Law}(Z_{2,t})$ both coincide with the original mean-field law $\mu_{t}$ for all $t\in[0,T]$. In particular, if $Z_{1,0}$ and $Z_{2,0}$ are identically distributed (with law $\mu_{0}$), then $Z_{1,t}\stackrel{{\scriptstyle d}}{{=}}Z_{2,t}$ for all $t\in[0,T]$, and this common distribution evolves according to the McKean--Vlasov equation (14) (proof in Appendix A.3). Thus, $\pi_{t}=\textrm{Law}(Z_{t})$ can be seen as the coupling of two independent copies of the mean-field solution.
Next, we consider the natural particle approximation of the two-system SDE. We introduce a coupled $2N$-particle system:
$\begin{split}dX_{t}^{j}\ &=\ b(t,X_{t}^{j},\delta_{\mathbf{Y}_{t}})\,dt\ +\ \sigma(t,X_{t}^{j},\delta_{\mathbf{Y}_{t}})\,dW_{t}^{X,j},\\ dY_{t}^{j}\ &=\ b(t,Y_{t}^{j},\delta_{\mathbf{X}_{t}})\,dt\ +\ \sigma(t,Y_{t}^{j},\delta_{\mathbf{X}_{t}})\,dW_{t}^{Y,j},\qquad j\in\{1,\ldots,N\},\end{split}$ (20)
where $\mathbf{X}_{t}=\{X_{t}^{j}\}_{j=1}^{N}$ and $\mathbf{Y}_{t}=\{Y_{t}^{j}\}_{j=1}^{N}$. Each of the $2N$ processes sees drift and diffusion terms driven by the empirical distribution of the opposite subsystem. By construction, (20) reduces to two independent copies of the original interacting particle system in the limit $N\to\infty$ (Appendix A.4).
Remark 1.
Our method shares similarities with conditional propagation of chaos [7], which utilizes the coupling
$\begin{split}dY^{0}_{t}\ &=\ b_{0}(t,Y^{0}_{t},\delta_{\mathbf{X}_{t}})\,dt\ +\ \sigma_{0}(t,Y^{0}_{t},\delta_{\mathbf{X}_{t}})\,dW^{0}_{t},\\ dX^{i}_{t}\ &=\ b(t,X^{i}_{t},\delta_{\mathbf{X}_{t}},Y^{0}_{t})\,dt\ +\ \sigma(t,X^{i}_{t},\delta_{\mathbf{X}_{t}},Y^{0}_{t})\,dW^{i}_{t},\quad\text{for}\quad i\in\{1,\ldots,N\},\end{split}$ (21)
and $\delta_{\mathbf{X}_{t}}=(\sum_{i=1}^{N}\delta_{X^{i}_{t}})/N$. This scheme is often utilized in maximum likelihood estimation [18, 20].
These methods use $\mathbf{X}_{t}$ to approximate an integral in the dynamics of a different system, $Y^{0}_{t}$, making $\textrm{Law}(Y^{0}_{t})$ a distribution conditioned on $\mathbf{X}_{t}$.
Our approach is structurally similar but differs in key respects. In the conditional propagation of chaos setup, $\mathbf{X}_{t}$ also determines the dynamics of its own particles, introducing correlations between them. In contrast, our two-system method uses the empirical law of $\mathbf{Y}_{t}$ to drive the dynamics of $\mathbf{X}_{t}$, and vice versa. As a result, conditioned on $\mathbf{Y}_{t}$, $X_{t}^{j}$ are independent, and vice versa. This conditional independence reduces intra-system correlation and more closely resembles ensemble-chain strategies used in discrete-time sampling.
5 A Two-System Approach for Samplers
For the rest of this paper, unless otherwise stated, we consider a restricted class of McKean--Vlasov equations
$d\overline{X}_{t}\ =\ C(\mu_{t})b(\overline{X}_{t})\,dt\ +\ \sqrt{2C(\mu_{t})}\,dW_{t},\qquad t\in[0,T],$ (22)
where $\mu_{t}=\textrm{Law}(\overline{X}_{t})$, and $C(\mu)$ is Lipschitz continuous, bounded, and positive definite for all $\mu\in\mathcal{P}(\mathbb{R}^{d})$. In particular, $b(\overline{X}_{t})$ no longer depends on $t$ and the law of $\overline{X}_{t}$. As such, for $b(x)=-\nabla V(x)$, the invariant distribution takes the form $\rho(x)\propto\exp(-V(x))$. Thus, the two-system framework introduces a pair of interacting Markov chains, where each chain generates proposals based on the empirical law of the other. When applied to a continuous-time mean-field sampler, this coupling preserves $\rho$ as the invariant distribution in both the finite-particle setting and the infinite-particle limit. By adding Metropolis--Hastings acceptance steps to an alternating (staggered) time-discretization of this coupled system, we obtain an ensemble-chain sampler that exactly maintains the target distribution $\rho$.
This connection demonstrates that ensemble-chain samplers approximate expectations with respect to $\rho$ by leveraging particle interactions. In the limit as $N\to\infty$ (and as time $t\to\infty$), the ensemble statistics converge to their mean-field counterparts. This also motivates the use of adaptive samplers, which replace the whole ensemble with a single chain and rely on long-run time averages to recover the same mean-field behavior.
5.1 Continuous-Time Samplers
To illustrate, consider the mean-field overdamped Langevin example (11). The two-system construction leads to a coupled SDE on $Z_{t}=\{X_{t},Y_{t}\}$ (with $X_{t},Y_{t}\in\mathbb{R}^{d}$):
$\begin{split}dX_{t}\ &=\ C(p_{1}\!\circ\mu_{t})\,\nabla\log\rho(X_{t})\,dt\;+\;\sqrt{2\,C(p_{1}\!\circ\mu_{t})}\,dW_{t}^{X},\\ dY_{t}\ &=\ C(p_{2}\!\circ\mu_{t})\,\nabla\log\rho(Y_{t})\,dt\;+\;\sqrt{2\,C(p_{2}\!\circ\mu_{t})}\,dW_{t}^{Y},\end{split}$ (23)
for $t\in[0,T]$, where $\mu_{t}=\textrm{Law}(Z_{t})$, and $\rho(X_{t})$
Under mild regularity conditions on $\nabla\log\rho$, the two-system McKean--Vlasov equation (23) admits a unique strong solution (e.g., if $\log\rho(x)$ is strictly concave outside a large ball).
Applying the finite-particle approximation from the previous section, we obtain the following $2N$-particle system for the two coupled overdamped processes:
$\begin{split}dX^{j}_{t}\ &=\ C(\delta_{\mathbf{Y}_{t}})\,\nabla\log\rho(X^{j}_{t})\,dt\;+\;\sqrt{2\,C(\delta_{\mathbf{Y}_{t}})}\,dW^{X,j}_{t},\\ dY^{j}_{t}\ &=\ C(\delta_{\mathbf{X}_{t}})\,\nabla\log\rho(Y^{j}_{t})\,dt\;+\;\sqrt{2\,C(\delta_{\mathbf{X}_{t}})}\,dW^{Y,j}_{t},\end{split}$ (24)
for $j=1,\dots,N$. By the result of Section 4, as $N\to\infty$, this system converges to the mean-field dynamics (23). Crucially, the two-system coupling also ensures exact invariance of the target at finite $N$. In particular, as we prove below (Lemma 1), $\rho^{\otimes 2N}$ is an invariant density of the coupled $2N$-particle system (24) for any $N$. In other words, even before taking $N\to\infty$, the two-system construction preserves the desired target distribution by design.
Lemma 1.
The $2N$-particle system (24) has $\rho^{\otimes 2N}$ as its invariant density.
Proof.
Using $Z_{t}=(Z_{1,t},Z_{2,t})$ where $Z_{1,t}=\mathbf{X}_{t}$ and $Z_{2,t}=\mathbf{Y}_{t}$, (24) can be written as
$dZ_{t}\ =\ S(Z_{t})\,\nabla\log\{\rho(Z_{t})\}\ +\ \sqrt{2S(Z_{t})}\,dW_{t},\qquad\text{where}\qquad S(Z_{t})\ =\ \begin{pmatrix}C(Z_{2,t})&0\\ 0&C(Z_{1,t})\end{pmatrix}$. (25)
Define
$(\nabla\cdot S(Z_{t}))_{i}\ =\ \sum_{j=1}^{2N}\partial_{j}S(Z_{t})_{ij},\qquad\text{for}\qquad i\in\{1,\ldots,2N\}.$ (26)
Then $\nabla\cdot S(Z_{t})=0$ is a sufficient condition for the corresponding Fokker–Planck to have $\rho^{\otimes 2N}$ as the invariant density [32, 34]. Since $S(Z_{t})_{i,j}$ does not depend on $(Z_{t})_{j}$, $(\nabla\cdot S(Z_{t}))_{i}=0$ for all $i\in\{1,\ldots,2N\}$. Thus $\rho^{\otimes 2N}$ is the invariant density of (24). ∎
This result shows that the two-system approach provides a principled way to transform a continuous-time mean-field sampler into a finite-particle sampler while maintaining $\rho$ as the invariant measure for any ensemble size.
Remark 2.
Lemma 1 can be extended to the case where the McKean–Vlasov equation takes the form
$d\overline{X}_{t}\ =\ C(\overline{\mu}_{t})b(t,\overline{X}_{t},\overline{\mu}_{t})\,dt\ +\ \sqrt{2C(\overline{\mu}_{t})}\,\sigma(t,\overline{X}_{t},\overline{\mu}_{t})\,dW_{t},$ (27)
and $\overline{\mu}_{t}=\textrm{Law}(\overline{X}_{t})$. In this case, the coupled SDE’s look like
$d\overline{X}_{t}\$$=\ C(\overline{\rho}_{t})b(t,\overline{X}_{t},\overline{\mu}_{t})\,dt\ +\ \sqrt{2C(\overline{\rho}_{t})}\,\sigma(t,\overline{X}_{t},\overline{\mu}_{t})\,dW_{t},$ (28)
$d\overline{Y_{t}}\$$=\ C(\overline{\mu}_{t})b(t,\overline{Y}_{t},\overline{\rho}_{t})\,dt\ +\ \sqrt{2C(\overline{\mu}_{t})}\,\sigma(t,\overline{Y}_{t},\overline{\rho}_{t})\,dW_{t},$ (29)
where $\overline{\rho}_{t}=\textrm{Law}(\overline{Y}_{t})$. As such, $\overline{Y}_{t}$ is only used to precondition the dynamics of $\overline{X}_{t}$, but is not used to compute $b$ or $\sigma$. As in Lemma 1, this satisfies $\nabla\cdot S(Z_{t})=0$, thus preserving the invariant density $\rho^{\otimes 2}$.
To extend the construction to the underdamped Langevin (10), introduce two coupled copies $(X_{t},V_{t})$ and $(Y_{t},U_{t})$ and denote
$\Pi_{t}=\textrm{Law}(X_{t},V_{t},Y_{t},U_{t}),\qquad\mu_{t}^{X}=p_{X}\circ\Pi_{t},\quad\mu_{t}^{Y}=p_{Y}\circ\Pi_{t}$
as the position marginals. The two-system kinetic Langevin is
$\begin{split}dV_{t}&=-\alpha V_{t}\,dt\;+\;\gamma\,C(\mu_{t}^{Y})^{1/2}\nabla\log\rho(X_{t})\,dt\;+\;\sqrt{2\alpha\gamma}\,dW_{t}^{V},\\ dX_{t}&=C(\mu_{t}^{Y})^{1/2}V_{t}\,dt,\\ dU_{t}&=-\alpha U_{t}\,dt\;+\;\gamma\,C(\mu_{t}^{X})^{1/2}\nabla\log\rho(Y_{t})\,dt\;+\;\sqrt{2\alpha\gamma}\,dW_{t}^{U},\\ dY_{t}&=C(\mu_{t}^{X})^{1/2}U_{t}\,dt,\end{split}$ (30)
again with cross-preconditioning via the other subsystem’s position law. When $C(\cdot)\equiv C_{0}$ is constant, (30) reduces to two independent preconditioned kinetic Langevin processes. The corresponding finite particle discretization and its convergence to $\rho$ naturally follow.
5.2 Discrete-Time Samplers
To obtain a practical MCMC algorithm, we discretize the two-system dynamics in time using an alternating (staggered) update scheme. Specifically, consider a time step $h>0$ and suppose we have the coupled system state $(X_{t},Y_{t})$ at time $t$. A single two-system update (of duration $h$) can be defined as:
$\begin{split}X_{t+h}\ &=\ X_{t}\;+\;h\,C(\mu_{t}^{Y})\,\nabla\log\rho(X_{t})\;+\;\sqrt{2h\,C(\mu_{t}^{Y})}\,\xi^{X},\\ Y_{t+h}\ &=\ Y_{t}\;+\;h\,C(\mu_{t+h}^{X})\,\nabla\log\rho(Y_{t})\;+\;\sqrt{2h\,C(\mu_{t+h}^{X})}\,\xi^{Y},\end{split}$ (31)
where $\mu_{t}^{Y}$ denotes the empirical law of subsystem $Y$ at time $t$, and $\xi^{X},\xi^{Y}$ are independent standard normal increments. In other words, we first update $X$ using the current $Y$-statistics, then update $Y$ using the new $X$-statistics (this is a symmetric split step akin to a leapfrog). This yields a sequence of proposals $(X_{t},Y_{t})\to(X_{t+h},Y_{t})\to(X_{t+h},Y_{t+h})$ for the coupled system. We then apply a Metropolis--Hastings accept/reject step at each proposal to ensure detailed balance with respect to $\rho$.
Concretely, in the finite-particle setting, this becomes an ensemble-chain sampler, where the state space $\mathbf{Z}_{t}=\{\mathbf{X}_{t},\mathbf{Y}_{t}\}$ evolves with alternating updates
$\mathbf{X}_{t+1}\ \sim\ \prod_{i=1}^{N}P_{\delta_{\mathbf{Y}_{t}}}(X_{t}^{i},\cdot),\qquad\text{with}\qquad\mathbf{Y}_{t+1}\ \sim\ \prod_{i=1}^{N}P_{\delta_{\mathbf{X}_{t+1}}}(Y_{t}^{i},\cdot),$ (32)
yielding $\mathbf{Z}_{t+1}=\{\mathbf{X}_{t+1},\mathbf{Y}_{t+1}\}$. This is exactly an ensemble chain MCMC method, where one group of particles proposes using the empirical distribution of the other group. Since each transition satisfies detailed balance with respect to $\rho$, the full scheme preserves $\rho^{\otimes 2N}$ (see Appendix A.6 for a proof). Then, in the mean-field limit, $N\to\infty$, the two-system variant of discrete-time mean-field samplers (2) emerges. In essence, this is a systematic way to obtain leave-one-out ensemble samplers without having to take step sizes to zero [41] or incur $O(N)$ costs per iteration, while maintaining the correct invariant distribution.
5.3 Adaptive Samplers
In the mean-field Langevin examples above, if $\Theta(\mu_{t})$ represents the covariance of $\mu_{t}$, then as $t\to\infty$ the mean-field dynamics $\Theta(\mu_{t})$ tend to the true covariance of $\rho$. Adaptive samplers mimic this behavior using a single chain in the long-time limit. One can view adaptive algorithms as approximating the mean-field ensemble behavior by replacing spatial averaging with temporal averaging.
Notably, ensemble methods themselves can incorporate adaptation. Given an ensemble $\mathbf{X}_{t}$, we could update an adaptive parameter via
$\Theta_{T}\ =\ \frac{1}{TN}\sum_{i=1}^{N}\theta(X_{T}^{i})\ +\ \Big(1-\frac{1}{T}\Big)\Theta_{T-1}\,$, (33)
for some statistic $\theta$. This effectively computes an average using all $N$ chains over time $T$. The corresponding ensemble transition kernel $\mathbf{P}_{\Theta_{T}}$ (which factorizes over $N$ independent chain updates with parameter $\Theta_{T}$) remains geometrically ergodic under mild conditions (by a Cauchy-Schwarz argument).
In a two-system adaptive variant, we can separate the adaptation for each subsystem. For example, we could maintain separate estimates for the $X$ and $Y$ groups:
$\begin{split}\Theta_{T}^{X}&=\frac{1}{TN}\sum_{i=1}^{N}\theta(X_{T}^{i})\;+\;\Big(1-\frac{1}{T}\Big)\Theta_{T-1}^{X},\\ \Theta_{T}^{Y}&=\frac{1}{TN}\sum_{i=1}^{N}\theta(Y_{T}^{i})\;+\;\Big(1-\frac{1}{T}\Big)\Theta_{T-1}^{Y},\end{split}$ (34)
and then use $\Theta_{T}^{X}$ when proposing updates for subsystem $Y$, and $\Theta_{T}^{Y}$ when updating subsystem $X$. In this way, each half of the ensemble adapts based on the other half, maintaining cross-system independence at each step.
Remark 3.
Analyzing adaptive MCMC typically requires verifying two ingredients, geometric ergodicity of the frozen kernels and diminishing adaptation, which can be technically involved. An alternative is to use a finite adaptation schedule: choose a cutoff time $T$, update the tuning parameter $\Theta_{t}$ only for $t\leq T$, and then fix $\Theta=\Theta_{T}$ for $t>T$. After $T$, the algorithm reduces to a standard (non-adaptive) MCMC with kernel $P_{\Theta_{T}}$, so classical convergence results apply. In particular, if $P_{\Theta_{T}}$ is $\rho$-irreducible and aperiodic, the post-$T$ segment converges to the target and (after a short additional burn-in) can be treated as draws from a stationary chain.
The samplers we consider in Section 6 can be shown to be geometrically ergodic and satisfy diminishing adaptation. As such, we do not consider finite adaptive samplers in this paper.
6 Algorithmic Realizations of Two-System Samplers
We now describe concrete two-system MCMC algorithms derived from Langevin dynamics. Our focus is on two specific instances: an overdamped sampler based on MALA [2, 33], and an underdamped sampler based on MAKLA [3]. We emphasize, however, that the two-system approach is general and can be applied to other MCMC proposal families as well (see, e.g., [41] for alternative ensemble samplers).
For the ensemble-chain (multi-chain) setting, convergence to the target $\rho$ is ensured as long as each Metropolis--Hastings update satisfies detailed balance. We therefore present the two-system ensemble versions of MALA and MAKLA in Algorithms 1 and 2.
Adaptive two-system samplers require additional care in analysis, as their convergence depends on both the target distribution’s geometry and the adaptation scheme’s behavior. Below, we list sufficient conditions on $\rho$ that ensure convergence of our adaptive algorithms (proofs deferred to Appendix A.5). Algorithm 3 provides the adaptive two-system MALA, and Algorithm 4 gives the adaptive two-system MAKLA. The analogous one-system versions follow naturally by collapsing the two systems into one, so we omit those for brevity.
6.1 Assumptions for Adaptive MALA and MAKLA
Assumption 3.
The target distribution can be written in either of the forms $\rho(x)=h(x)e^{-p(x)}$ or $\rho(x)=h(x)^{-p(x)}$, where $h(x)\geq 0$ and $h(\cdot),p(\cdot)$ are polynomial functions.
Assumption 4.
The target distribution satisfies: (i) $\log\rho(x)$ is strongly concave (log-strongly-concave target), (ii) $\nabla\log\rho(x)$ is Lipschitz continuous, and (iii) the third derivatives of $\log\rho(x)$ are uniformly bounded in a subquadratic sense, namely there exists $L_{H}$ such that
$\Big|\sum_{i,j,k}\partial_{ijk}^{3}\log\rho(x)\;a_{i}b_{j}c_{k}\Big|\ \leq\ L_{H}\,\|a\|\,\|b\|\,\|c\|,\qquad\forall\,x,a,b,c\in\mathbb{R}^{d}.$
Under Assumption 3, the adaptive truncated MALA algorithm (Algorithm 3) is known to be geometrically ergodic and to satisfy diminishing adaptation, following results of Atchadé [1] (proved in Appendix A.5). For MAKLA, Bou-Rabee and Oberdörster [3] show that geometric ergodicity holds under Assumption 4. Furthermore, the adaptive two-system MAKLA scheme (Algorithm 4) satisfies diminishing adaptation, since the adaptation magnitude decays on the order of $1/n$ and the mapping from the space of positive-definite matrices (with operator norm bounded by some constant $K$) to the proposal kernel is continuous.
We note that Riou-Durand et al. [36] recently proposed a finite adaptive scheme based on the MALT sampler. [^1] The MALT sampler is a Metropolized integrator on an OBABO discretization scheme for the underdamped Langevin diffusion, using full momentum refreshment at each step [37]. The MAKLA sampler utilizes the OABAO scheme, while keeping the momentum component and flipping the sign upon rejection. Our adaptive two-system MAKLA, in contrast, uses an infinite adaptation: the adaptation continues to evolve throughout the entire run rather than freezing after a burn-in phase. In practice, this leads to better long-term efficiency and robustness across a wider range of targets, especially when the target’s geometry is highly anisotropic or has spatially varying curvature.
Another distinction is that our adaptation explicitly preconditions the gradient $\nabla\log\rho(x)$ in the proposal step, whereas the adaptation in [37] preconditions the distribution used for momentum refreshment. By directly preconditioning the gradient updates (Section 6.4), our algorithm aligns the dynamics with the local curvature of $\log\rho(x)$, accelerating exploration along stiff directions and mitigating random-walk behavior. This is particularly important in high dimensions, where poorly scaled gradients can severely slow convergence and reduce effective sample size.
Algorithm 1 Ensemble Chain MALA
0: Desired distribution $\rho:\mathbb{R}^{d}\to[0,\infty)$, starting points $(\mathbf{X}_{0}^{s})_{j}\in\mathbb{R}^{d}$ for $s=0,1$, and $j\in\{1,\ldots,N\}$, number of MCMC samples $M\geq 1$, step size $h>0$, and $e\geq 0$ to ensure positive-definite covariance
$s\leftarrow 0$
for $m\leftarrow 1$, $m\leq M$ do
  $C^{1-s}_{m}\leftarrow C(\mathbf{X}_{m}^{1-s})+eI_{d}\qquad$ (Refresh covariance)
  (Accept or reject particles individually)
  for $j\in\{1,\ldots,N\}$ (in parallel) do
   $(\mathbf{X}_{m}^{s})_{j}\leftarrow\text{ProposeMALA}_{h}((\mathbf{X}_{m-1}^{s})_{j},{C^{1-s}_{m}})$
  end for
  $s\leftarrow 1-s\qquad$ (Switch systems)
end for
Procedure $\textrm{ProposeMALA}_{h}(X_{0},C)$
$\quad X_{1}\sim\mathcal{N}(X_{0}+hC\nabla\log\rho(X_{0}),2hC)\qquad$ (Propose a Langevin trajectory)
$\quad\alpha\leftarrow\textrm{AcceptRatioMALA}(X_{0},X_{1},C)$
$\quad\beta\sim U(0,1)$
  if $\beta>\alpha$ then
   return $X_{0}$   (Reject)
  end if
  return $X_{1}$   (Accept)
end Procedure
Procedure $\textrm{AcceptRatioMALA}(X,Y,C)$
$\quad\alpha\leftarrow\frac{\rho(Y)Q_{C}(X\mid Y)}{\rho(X)Q_{C}(Y\mid X)},\qquad\text{where}\ Q_{C}(y\mid x)\ \textrm{is the p.d.f. of $\mathcal{N}(x+hC\nabla\log\rho(x),2hC)$}$
$\quad\textbf{return}\ \min(1,\alpha)$
end Procedure
Algorithm 2 Ensemble Chain MAKLA
0: Desired distribution $\rho:\mathbb{R}^{d}\to[0,\infty)$, starting points $(\mathbf{X}_{0}^{s})_{j},(\mathbf{V}_{0}^{s})_{j}\in\mathbb{R}^{d}$ for $s=0,1$, and $j\in\{1,\ldots,N\}$, number of MCMC samples $M\geq 1$, step size $h>0$, leap-frog steps $L>0$, persistence $\eta\in(0,1)$, and $e\geq 0$ to ensure positive-definite covariance
$s\leftarrow 0$
for $m\leftarrow 1$, $m\leq M$ do
  $C^{1-s}_{m}\leftarrow C(\mathbf{X}_{m}^{1-s})+eI_{d}\qquad$ (Refresh covariance)
  (Accept or reject particles individually)
  for $j\in\{1,\ldots,N\}$ (in parrallel) do
   $((\mathbf{X}_{m}^{s})_{j},(\mathbf{V}_{m}^{s})_{j})\leftarrow\text{ProposeMAKLA}^{L}_{h,\eta}((\mathbf{X}_{m-1}^{s})_{j},(\mathbf{V}_{m-1}^{s})_{j},{C^{1-s}_{m}})$
  end for
  $s\leftarrow 1-s\qquad$ (Switch systems)
end for
Procedure $\textrm{ProposeMAKLA}^{L}_{h,\eta}(X_{0},V_{0},C)$
$\quad(X_{L},V_{L},\Delta)\leftarrow\textrm{OABAO}^{L}_{h,\eta}(X_{0},V_{0},C)\qquad$ (Propose a Langevin trajectory)
$\quad A\sim U(0,1)$
  if $A>\exp(-\Delta)$ then
   return $(X_{0},-V_{0})$   (Reject and flip momentum)
  end if
  return $(X_{L},V_{L})$   (Accept)
end Procedure
Procedure $\textrm{OABAO}^{L}_{h,\eta}(X_{0},V_{0},C)$
$\quad\Delta\leftarrow 0$
   for i=0;i<Li=0;\,i<L do
    (O) $V_{i}^{\prime}\leftarrow\sqrt{1-\eta}V_{i}+\sqrt{\eta}\xi_{i}^{1}$.
    (A) $X_{i+1/2}\leftarrow X_{i}+\frac{h}{2}\,C^{1/2}V_{i}^{\prime}$.
    (B) $V_{i+1}^{\prime}\leftarrow V_{i}^{\prime}+hC^{1/2}\nabla\log\rho(X_{i+1/2})$
    (A) $X_{i+1}\leftarrow X_{i+1/2}+\frac{h}{2}\,C^{1/2}V_{i+1}^{\prime}$.
    (O) $V_{i+1}\leftarrow\sqrt{1-\eta}V_{i+1}^{\prime}+\sqrt{\eta}\xi_{i}^{2}$
$\quad\quad\Delta\leftarrow\Delta+H(X_{i},V_{i}^{\prime})-H(X_{i+1},V^{\prime}_{i+1})\qquad$ (Gather local Hamiltonian errors)
  end for
  return $(X_{L},V_{L},\Delta)$
end Procedure
Algorithm 3 Two System Adaptive Ensemble MALA
0: Desired distribution $\rho:\mathbb{R}^{d}\to[0,\infty)$, starting points $(\mathbf{X}_{0}^{s})_{j}\in\mathbb{R}^{d}$ for $j\in\{1,\ldots,N\}$ and $s=0,1$, number of MCMC samples $M\geq 1$, step size $h>0$, drift truncation $\delta>0$, covariance truncation $K>0$, and small $\varepsilon>0$
$K\leftarrow 1$   (Adaptive counter)
for $m\leftarrow 1$, $m\leq M$ do
  for $s\in\{0,1\}$ do
   $C^{1-s}_{m}\leftarrow C^{1-s}(\mathbf{X}_{m}^{s})/K+(1-1/K)C^{1-s}_{m-1}\qquad$ (Adapt covariance)
   $C_{m}^{1-s}\leftarrow\frac{K}{\max(K,\lvert C^{1-s}_{m}\rvert)}C^{1-s}_{m}$
   (Accept or reject particles)
   for $j\in\{1,\ldots,N\}$ (in parallel) do
    $(\mathbf{X}^{s}_{m})_{j}\leftarrow\text{ProposeTruncatedMALA}_{h,\delta}((\mathbf{X}_{m-1}^{s})_{j},{C_{m}^{1-s}+\varepsilon I_{d}})$
   end for
   $K\leftarrow K+1$
  end for
end for
Procedure $\textrm{ProposeTruncatedMALA}_{h,\delta}(X,C)$
$\quad d\leftarrow D(X),\qquad$ where $D(x)$ defined in (107)
$\quad Y\sim\mathcal{N}(X+hCd,2hC)\qquad$ (Propose a Langevin trajectory)
$\quad\alpha\leftarrow\textrm{AcceptRatioTruncatedMALA}(X,Y,C)$
$\quad\beta\sim U(0,1)$
  if $\beta>\alpha$ then
   return $X$   (Reject)
  end if
  return $Y$   (Accept)
end Procedure
Procedure $\textrm{AcceptRatioTruncatedMALA}_{h,\delta}(X,Y,C)$
$\quad\alpha\leftarrow\frac{\rho(Y)Q_{C}(X\mid Y)}{\rho(X)Q_{C}(Y\mid X)},\qquad\text{where}\ Q_{C}(y\mid x)\ \textrm{is the p.d.f. of $\mathcal{N}(x+hCD(x),2hC)$}$
$\quad\textbf{return}\ \min(1,\alpha)$
end Procedure
Algorithm 4 Two System Adaptive Ensemble MAKLA
0: Desired distribution $\rho:\mathbb{R}^{d}\to[0,\infty)$, starting points $(\mathbf{X}^{s}_{0})_{j},(\mathbf{V}^{s}_{0})_{j}\in\mathbb{R}^{d}$ for $j\in\{1,\ldots,N\}$ and $s\in\{0,1\}$, number of MCMC samples $M\geq 1$, step size $h>0$, leap-frog steps $L>0$, persistence $\eta\in(0,1)$, and small $\varepsilon>0$
$K\leftarrow 1$   (Adaptive counter)
for $m\leftarrow 1$, $m\leq M$ do
  for $s\in\{0,1\}$ do
   $C^{1-s}_{m}\leftarrow C(\mathbf{X}^{1-s}_{m})/K+(1-1/K)C^{1-s}_{m-1}\qquad$ (Adapt covariance)
   $C^{1-s}_{m}\leftarrow\frac{K}{\max(K,\lvert C^{1-s}_{m}\rvert)}C^{1-s}_{m}$
   (Accept or reject particles)
   for $j\in\{1,\ldots,N\}$ (in parallel) do
    $((\mathbf{X}^{s}_{m})_{j},(\mathbf{V}^{s}_{m})_{j})\leftarrow\text{ProposeMAKLA}^{L}_{h,\eta}((\mathbf{X}^{s}_{m-1})_{j},{C^{1-s}_{m}+\varepsilon I_{d}})$
   end for
   $K\leftarrow K+1$
  end for
end for
6.2 Step Size Control via Randomization
The performance of Langevin-based methods is highly sensitive to the choice of step size $h$. Large values of $h$ reduce autocorrelation in the resulting Markov chain, but if $h$ is too large, proposals are frequently rejected, again leading to poor mixing.
A rejection can be interpreted as evidence that the chain is locally sensitive to $\mathcal{O}(h)$ perturbations, and hence that the discretization error of the underlying SDE is too large. A classical remedy is delayed rejection [31, 43], in which a rejected proposal is followed by one or more subsequent proposals with progressively smaller step sizes. While theoretically attractive, this approach has two key drawbacks: (i) the computational cost of the $k^{\text{th}}$ proposal scales as $\mathcal{O}(2^{k})$, and (ii) it is poorly suited to modern CPU/GPU hardware, where Single Instruction Multiple Data (SIMD) parallelism requires all chains to follow the same instruction path.
As an alternative, we introduce a randomized step size
$h=\gamma h_{\max},\qquad\gamma\sim\beta\,\delta_{x-1}+(1-\beta)\,f(x),$ (35)
where $f(x)$ is a distribution on $[0,1]$ with no mass at zero, and $0<\beta<1$, pictorially demonstrated in Fig. 1. With probability $\beta$, the maximum step size $h_{\max}$ is used; otherwise, a random fraction $c\in(0,1)$ is drawn from $f(x)$, and $ch_{\max}$ is used. This randomized scheme preserves reversibility and can be applied independently or jointly across chains without disturbing invariance under $\rho^{\otimes 2N}$. Crucially, it automatically introduces occasional smaller step sizes when proposals are frequently rejected, thereby stabilizing the dynamics.
The choice of $f$ governs how often smaller steps are taken. A uniform distribution on $[0,1]$ is natural, but does not sufficiently emphasize small step sizes. We therefore adopt a distribution with density $f(x)=3(1-x)^{2}$ on $[0,1]$, which biases the sampler toward smaller $h$ values while still allowing exploration with larger steps. The expected step size is then
$\mathbb{E}[h]=h_{\max}\int_{0}^{1}x\big(\beta\delta(x-1)+3(1-\beta)(1-x)^{2}\big)\,dx=h_{\max}\left(\beta+\frac{1-\beta}{4}\right).$ (36)
This randomized step size scheme is particularly effective on funnel-like targets (e.g., Neal’s funnel), where strong anisotropy creates regions of large curvature that demand small steps, interspersed with regions of low curvature where larger steps are essential for efficient exploration.
Figure 1: Visualization of the randomized step size distribution. The step size $h=\gamma h_{\max}$ is drawn from a mixture of a point mass at $\gamma=1$ with weight $\beta\in(0,1)$, and a continuous component $f(x)=3(1-x)^{2}$ supported on $(0,1)$, with weight $1-\beta$. This construction encourages frequent large proposals while allowing occasional small, exploratory steps, improving robustness across varying curvature scales.
6.3 Restarting Adaptive Schemes
The adaptive updates we consider in this paper are implemented via a running estimator
$\Theta_{K}\;=\;\left(1-\frac{1}{K}\right)\,\Theta_{K-1}\;+\;\frac{1}{K}\,\theta(X_{K}),$
for some statistic $\theta(X_{K})$ (covariance, step size, etc.). Since the iteration $K$ increases, later samples have diminishing influence on $\Theta_{K}$; yet the early samples (drawn far from stationarity) receive disproportionate weight. This mismatch slows the convergence of $\Theta_{K}$ to $\mathbb{E}_{\rho}[\theta(X)]$ and creates a negative feedback loop: poor $\Theta_{K}\Rightarrow$ poor proposals $\Rightarrow$ slow mixing $\Rightarrow$ poor $\Theta_{K}$. With ensembles sharing a common $\Theta_{K}$, the problem is exacerbated by sensitivity to initialization.
To mitigate this, during a burn-in of length $T_{\mathrm{burn}}$, we restart the adaptation on a fixed schedule
$\mathcal{R}\;=\;\{\tau,\,2\tau,\,\dots,\,\tau_{\max}\},\quad\text{with }\tau_{\max}<T_{\mathrm{burn}}.$
At each $t\in\mathcal{R}$ we either (i) hard-reset the estimator (reset any counters controlling $\Theta_{K}$ and forget any information from the historical samples, e.g., setting $K=1$), or (ii) soft-reset the estimator (keep $\Theta_{K}$ but reset the schedule so that future samplers have a larger impact, e.g., setting K=2K=2). [^2] In both cases, $K$ no longer corresponds to the number of samples generated by the sampler. No restarts are performed after $\tau_{\max}$; we then allow a short stabilization window (the samples between $\tau_{\max}$ and $T_{\text{burn}}$) and begin collecting samples. In our experiments, we take $\tau_{\max}=0.5\,T_{\mathrm{burn}}$ and use hard resets, as we find that hard resets yield slightly higher acceptance ratios than soft resets. This method is summarized in Algorithm 5 for a general (one-system) ensemble adaptive sampler.
These restarts facilitate faster convergence to the stationary covariance estimate and reduce the size of the burn-in phase required for optimal efficiency. Practically, this leads to better-calibrated preconditioners, higher acceptance rates, and improved ESS per gradient during the sampling phase, while preserving the usual guarantees of adaptive MCMC samplers.
Algorithm 5 Restarted Ensemble Adaptive Sampler (shared statistic $\Theta$ with hard or soft resets)
0: Target $\rho$, proposal kernel family $Q_{\Theta}$, burn-in $T_{\mathrm{burn}}$, total iterations $T_{\mathrm{tot}}$, restart period $\tau$, last restart time $\tau_{\max}<T_{\mathrm{burn}}$
 Initialize chains $\{X_{0}^{(n)}\}_{n=1}^{N}$, statistic $\Theta\leftarrow\Theta_{\mathrm{init}}$
$K\leftarrow 1$ (Adaptation counter)
for $t\leftarrow 1$; $t\leq T_{\mathrm{tot}}$ do
  for $n\leftarrow 1$; $n\leq N$ (in parallel) do
   Propose $Y\sim Q_{\Theta}(\cdot\mid X_{t-1}^{(n)})$; accept with prob. $\alpha$; set $X_{t}^{(n)}\leftarrow Y$ or $X_{t-1}^{(n)}$
  end for
  $\Theta\leftarrow(1-1/K)\,\Theta\;+\;{\theta}(\{X_{t}^{(n)}\}_{n=1}^{N})/K$
  $K\leftarrow K+1$
  if $t\leq T_{\mathrm{burn}}$ and $t\in\{\tau,2\tau,\dots,\tau_{\max}\}$ then
   if HardReset then
    $K\leftarrow 1$   (Hard reset)
   else
    $K\leftarrow 2$   (Soft reset)
   end if
  end if
end for
6.4 Dealing with Ill-Conditioned Distributions
Posterior distributions arising in practice (e.g., from posteriordb) are often severely ill-conditioned. For example, the Hessian at the maximum a posteriori can have eigenvalues with a dynamic range as large as $10^{11}$. This creates an extremely anisotropic gradient field, where the log-density is very steep in some directions and nearly flat in others.
Momentum Refresh: For samplers that rely on integrating a chain of gradient evaluations within a single proposal (such as MALT, HMC, and MAKLA), multiple leapfrog steps can easily lead to numerical instability, often producing explosive trajectories with infinite or NaN values. Stability can be restored by dramatically shrinking the step size, but this in turn reduces sample efficiency. A more robust strategy is to use only a single leapfrog step, which permits a larger step size without sacrificing stability. We have adopted this strategy in all of our experiments.
However, in this regime, the way momentum is refreshed becomes crucial. HMC and MALT employ full velocity resampling at every proposal, which lowers efficiency by discarding information about the current momentum. For HMC, this resampling is essential to guarantee convergence to the invariant distribution (via the Virial theorem). In contrast, MALT and MAKLA approximate the underdamped Langevin dynamics, for which momentum refreshment is already built into the dynamics. Moreover, given any initial position and velocity, the continuous-time process converges to the correct invariant distribution, and the Metropolis correction merely removes discretization bias. Hence, additional momentum resampling is unnecessary, and retaining momentum (with sign reversal on rejection) yields more efficient sampling. For this reason, we adopt MAKLA as the base sampler.
Rescaling the Distribution: Ill-conditioning also implies that different coordinates of the state space live on vastly different scales. The smallest scales, associated with the largest curvature directions, force the step size $h$ to be excessively small. To mitigate this, we rescale the distribution as
$g(z_{1},\ldots,z_{d})\ =\ \rho(a_{1}z_{1},\ldots,a_{d}z_{d}),\qquad a_{i}>0,$ (37)
which differs from $\rho(x)$ only by a normalization constant. Thus, no Jacobian correction is required, and we may sample from $g(z)$ in place of $\rho(x)$. This rescaling balances the coordinate scales, enabling larger, stable step sizes and improving sampling efficiency while mitigating numerical instabilities. Finally, samples from the original distribution $\rho(x)$ are recovered by rescaling the draws from $g(z)$ via $x_{i}=a_{i}z_{i}$. For our experiments, we (numerically) find the maximum $x^{\star}$ of $\rho$ and compute the Hessian $H=-\nabla^{2}\log\rho(x^{\star})$. We then define a diagonal scaling as
$a_{i}\;=\;\frac{1}{\sqrt{\,H_{ii}+\varepsilon\,}},\qquad\text{for small}\qquad\varepsilon>0,$
so that $1/a_{i}=\sqrt{H_{ii}+\varepsilon}$ is the square root of the $i$-th diagonal entry of the negative log-density Hessian at $x^{\star}$ (with $\varepsilon$ for numerical stability). The resulting diagonal preconditioner is
$C_{0}\;=\;\operatorname{diag}(a_{1}^{2},\dots,a_{d}^{2})\;=\;\operatorname{diag}\!\big(1/(H_{11}+\varepsilon),\dots,1/(H_{dd}+\varepsilon)\big)\;=\;\big(\operatorname{diag}H+\varepsilon I\big)^{-1},$
i.e., a diagonal approximation to $(H+\varepsilon I)^{-1}$ that normalizes the local curvature near $x^{\star}$.
7 Experiments
We conducted a two-part empirical study. In Section 7.1, we evaluate the two-system samplers on controlled synthetic targets with known expectations (Gaussian, Student-$t$, and Neal’s funnel distributions) to measure bias and sampling efficiency. In Section 7.2, we move to real-data experiments, benchmarking our methods on 45 posterior distributions from posteriordb for which reference samples were provided, and comparing performance against NUTS.
In this section, to draw a distinction between ensemble-chain and adaptive-ensemble-chain samplers, we call a pure ensemble-chain MCMC method a ‘coupled’ sampler and an adaptive-ensemble-chain an ‘adaptive’ sampler. For the adaptive methods, we consider two adaptation modes: (i) 1sys-adaptive, which uses the full ensemble of $N$ chains to update a single shared covariance (as in (33)), and (ii) 2sys-adaptive, which splits the ensemble $N$ into two subsystems and adapts each half separately (as in (34)) to combine long-term adaptation with cross-chain independence.
7.1 Synthetic Experiments
To test the two-system samplers in scenarios where ground-truth is available, we consider distributions in $\mathbb{R}^{d}$ for which expectations can be computed in closed form. Specifically, we examine:
Neal’s Funnel (dim =10=10, $\sigma=3$): $\rho(x,y)\propto\mathcal{N}(x\mid 0,\sigma^{2})\prod_{i=1}^{d-1}\mathcal{N}(y_{i}\mid 0,\,e^{x}),$
Student-$t$ (dim =100=100, $\nu=4$): $\rho(x)\propto\Big(1+\frac{1}{\nu}\|A^{1/2}x\|^{2}\Big)^{-(d+\nu)/2},$
Gaussian (dim =100=100): $\rho(x)\propto\exp\!\Big(-\frac{1}{2}\|A^{1/2}x\|^{2}\Big),$
where for the Student-$t$ and Gaussian cases, $A$ is a randomly generated symmetric positive-definite matrix with eigenvalues linearly spaced in $10^{-2},10^{2}]$. These targets pose different challenges: Neal’s funnel is severely ill-conditioned (and often used to stress-test HMC and NUTS), the Student-$t$ is heavy-tailed, and the Gaussian is high-dimensional but benign in shape.
Experimental Setup. We compare standard MALA and MAKLA against their coupled and adaptive two-system variants on the above targets. The experiments in this section use $N=10$ parallel chains and $R=15$ independent repeated runs. In the 2sys-adaptive and coupled methods, each subsystem contains 5 chains. We run each sampler for 5 0005\,000 burn-in iterations followed by 15 00015\,000 sampling iterations, with one sample collected per iteration. Step sizes are tuned to target an acceptance rate around 0.7 for all samplers. For the adaptive schemes, we reset the adaptation parameters every 100 steps during burn-in (this “warm restarts” technique was critical for achieving high efficiency).
Performance Metrics. We evaluate sampler performance using two metrics: (i) ESS per gradient (ESS/Grad), which measures sampling efficiency per unit of work, and (ii) Gradient steps to low bias, which measures how quickly each method approaches the true distribution in terms of number of gradient evaluations:
  1. 1.
    ESS/Grad: For each coordinate $j=1,\dots,d$, we compute the effective sample size (ESS) from the final 1000 post-burn-in samples across all chains and runs. ESS for coordinate $j$ is estimated as
    $\text{ESS}_{j}=\frac{RNT\,\widehat{\mathrm{Var}}(x_{j})}{\widehat{\mathrm{Var}}(\bar{x}^{(r)}_{j})}\,$,
    where $RNT$ is the total number of samples ($15\text{ runs}\times 10\text{ chains}\times 1\,000$ samples), $\widehat{\mathrm{Var}}(x_{j})$ is the sample variance of all draws for coordinate $j$, and $\widehat{\mathrm{Var}}(\bar{x}^{(r)}_{j})$ is the variance of the run means (accounting for between-run variability). We report the median ESS across dimensions, divided by the total number of gradient evaluations performed (giving ESS per gradient computation).
  2. 2.
    Gradient Steps to Low Bias: Let $f:\mathbb{R}^{d}\to\mathbb{R}^{k}$ be a diagnostic vector function (e.g., low-order moments). Define the “bias” at iteration $t$ as
    $\mathcal{L}_{f,t}=\max_{1\leq i\leq k}\;\frac{\big(\hat{f}_{t,i}-\mathbb{E}[f_{i}(x)]\big)^{2}}{\mathrm{Var}[f_{i}(x)]}\,$,
    where $\hat{f}_{t,i}$ is the average of $f_{i}$ over all runs and chains at iteration $t$, and $\mathbb{E}[f_{i}(x)]$ and $\mathrm{Var}[f_{i}(x)]$ are the true values. The reported quantity is the number of gradient steps until $\mathcal{L}_{f,t}(\widehat{X})<0.01$. We use $f_{i}(x)=x_{i}$ for Neal’s funnel and $f(x)=\|xA^{1/2}\|$ for the Gaussian and Student-$t$ targets.
Results --- The ESS/Grad are shown in Table 1 and Table 2 for the Langevin and MAKLA samplers, respectively, with Table 3 and Table 4 showing the number of gradients required to achieve low bias for the Langevin and MAKLA samplers, respectively.
Adaptive vs Coupled Samplers: Across both MALA and MAKLA, adaptive variants typically outperform their coupled counterparts in ESS/Grad when a full covariance matrix is employed. For instance, in the Gaussian target, the full covariance 1sys-Adaptive MALA sampler achieves an ESS/Grad of $5.45\!\times\!10^{-2}$, an order of magnitude higher than the coupled version ($1.84\!\times\!0^{-3}$), and nearly $25$ fold higher than baseline MALA ($2.13\times 10^{-3}$). A similar trend is observed in the MAKLA setting, where full-covariance adaptive samplers (especially the two-system variant) attain ESS/Grad values of $0.118$ and $0.207$ for the Gaussian and StudentT targets, respectively—far surpassing both the diagonal counterparts and non-adaptive ensembles.
The performance of diagonal-covariance adaptive samplers is more muted. In MALA, for example, gains over the coupled version are marginal, and in some cases (e.g., Neal’s funnel), the diagonal 2sys-adaptive variant underperforms the simpler coupled sampler. This suggests that the diagonal parametrization limits the adaptive scheme’s ability to reshape the geometry effectively, especially in anisotropic or heavy-tailed settings.
Diagonal vs Full Covariance: The advantage of full covariance preconditioning is clearest in targets with correlated geometry. In the Gaussian and StudentT distributions, which possess non-trivial covariance structure, full preconditioning offers significant gains over diagonal parameterizations. Switching from diagonal to full preconditioning in adaptive MALA samplers yields up to a 25 fold improvement in ESS/Grad. In MAKLA, the effect is even more dramatic: for example, the 1sys-Adaptive MAKLA sampler improves from $2.78\times 10^{-3}$ to $0.173$ (StudentT), and from $2.77\times 10^{-3}$ to $0.113$ (Gaussian); an improvement of two orders of magnitude.
However, in Neal’s funnel distribution, each coordinate is conditionally independent given the scale, and the true posterior is axis-aligned. Consequently, diagonal and full covariance choices yield similar performance, and in some cases, diagonal preconditioning slightly outperforms full variants due to reduced estimation variance in high dimensions. This highlights that geometric adaptivity must align with the target structure to be effective.
MALA vs MAKLA: Across all benchmarks, MAKLA consistently outperforms MALA. For instance, the base MAKLA sampler achieves ESS/Grad of $8.64\!\times\!10^{-3}$ on the Gaussian target, compared to $2.13\!\times\!10^{-3}$ for MALA. Full-covariance adaptive variants amplify this advantage: 2sys-Adaptive MAKLA reaches $0.118$ (Gaussian) and $0.207$ (StudentT), compared to $0.0545$ and $0.0484$ in the MALA counterparts.
In terms of gradient efficiency to low bias, MAKLA also dominates. On the StudentT target, Full covariance 1sys-Adaptive MAKLA achieves low bias in just 1 1031\,103 gradients. On the Gaussian target, the Full covariance 2sys-Adaptive MAKLA sampler reaches the threshold in 1 8791\,879 gradients, outperforming all MALA variants. These results suggest that the underdamped dynamics improve long-term exploration and transient burn-in behavior compared to overdamped Langevin flows.
Table 1: Effective sample size per gradient evaluation (ESS/Grad) for MALA-based samplers across three target distributions. We compare vanilla MALA, coupled MALA, and one- and two-system adaptive variants, under both diagonal and full covariance parameterizations. Best (highest) values are highlighted in bold, worst (lowest) are underlined.
Diagonal Covariance Full Covariance
MALA Coupled 1sys-Adaptive 2sys-Adaptive Coupled 1sys-Adaptive 2sys-Adaptive
Neals 2.35e-03 2.38e-03 2.58e-03 3.37e-03 2.43e-03 3.72e-03 2.11e-03
StudentT 2.06e-03 2.04e-03 2.18e-03 2.13e-03 1.54e-03 4.84e-02 1.91e-02
Gaussian 2.13e-03 2.05e-03 2.18e-03 2.15e-03 1.84e-03 5.45e-02 4.88e-02
Table 2: Effective sample size per gradient evaluation (ESS/Grad) for MAKLA-based samplers across three target distributions. The table compares the baseline kinetic Langevin sampler, MAKLA, with coupled variants and adaptive two-system methods. Both diagonal and full covariance structures are considered. Best (highest) values are highlighted in bold, worst (lowest) are underlined.
Diagonal Covariance Full Covariance
MAKLA Coupled 1sys-Adaptive 2sys-Adaptive Coupled 1sys-Adaptive 2sys-Adaptive
Neals 4.12e-02 5.42e-03 1.13e-02 8.81e-03 4.12e-03 1.11e-02 1.05e-02
StudentT 4.76e-03 2.72e-03 2.78e-03 2.77e-03 2.27e-03 1.73e-01 2.07e-01
Gaussian 8.64e-03 2.58e-03 2.77e-03 2.81e-03 1.64e-03 1.13e-01 1.18e-01
Table 3: Number of gradient evaluations required for MALA-based samplers to reach low bias ($\mathcal{L}_{f,t}(\widehat{X})<0.01$). Samplers failing to reach this threshold are marked with a dash. The most gradient-efficient samplers are highlighted in bold; dashes indicate no convergence to the target bias level.
Diagonal Covariance Full Covariance
MALA Coupled 1sys-Adaptive 2sys-Adaptive Coupled 1sys-Adaptive 2sys-Adaptive
Neals 1911 10 395 2567 3051 - 6594 2190
StudentT - - - - - - 7179
Gaussian 5537 - 14 086 - - 5270 4322
Table 4: Number of gradient evaluations required for MAKLA-based samplers to reach low bias ($\mathcal{L}_{f,t}(\widehat{X})<0.01$). Samplers failing to reach this threshold are marked with a dash. The most gradient-efficient samplers are highlighted in bold; dashes indicate no convergence to the target bias level.
Diagonal Covariance Full Covariance
MAKLA Coupled 1sys-Adaptive 2sys-Adaptive Coupled 1sys-Adaptive 2sys-Adaptive
Neals 0 0 727 6 - 798 2 112
StudentT 7510 - - - - 1103 4167
Gaussian 3 1656 4561 711 - 1559 1879
7.2 Posteriordb Benchmarks: NUTS vs. Coupled MAKLA vs. Adaptive MAKLA
We benchmarked our adaptive and coupled MAKLA samplers against the No-U-Turn Sampler (NUTS) on 45 posterior distributions from posteriordb. posteriordb contained 47 examples with reference samples. However, we found that for two of them (eight_schools-eight_schools_centered and mcycle_gp-accel_gp), the posterior mode could not be found with any optimizer, and the norm of the gradient kept increasing during optimization, potentially due to an improper posterior. Hence, we have omitted these two examples from the comparison.
The coupled MAKLA sampler was run with $8d$ particles, where $d$ is the dimension of each posterior distribution, while the adaptive methods were run with $N=20$ parallel chains; for the two-system variants, this is split evenly between the two subsystems. The step size $h$ was pre-tuned to target an acceptance ratio of $\approx 1-h/4$, and each adaptive MAKLA system used $k=1$ OABAO step (we decreased $h$ starting from 1 by a factor of $\sqrt{2}$ at a time until acceptance rate exceeded 1h/41-h/4). For all MAKLA samplers, the velocity refreshment parameter $\eta$ was chosen as $\eta=\exp(-\gamma h)$ with $\gamma=1/16$. The number of total burn-in steps was chosen as $2\,000\cdot\lceil 1/h\rceil$, while the number of sampling steps was chosen as $4\,000\cdot\lceil 1/h\rceil$, with thinning of $\lceil 1/h\rceil$. All samplers were initialized at the mode of the distribution, which was found by the conjugate-gradient Newton method (Newton-CG from scipy). For fairness, we have also initiated NUTS at the mode of the distribution.
We report efficiency as the effective sample size per gradient evaluation (ESS/Grad). We consider two versions, the minimum ESS/Grad (minimum among all components), and median ESS/Grad (among all components). The ESS values and the Gelman-Rubin statistics were computed using Martyn Plummer’s coda package. Adaptation statistics were reset every $200\cdot\lceil 1/h\rceil$ proposals during burn-in (with resets stopped after 0.5 of the total burn-in steps). The posterior distributions are rescaled as outlined in Section 6.4.
Efficiency vs. dimension.
Figure 2 plots median ESS/Grad against dimension, with separate panels for NUTS, Coupled MAKLA, 1sys-Adaptive, and 2sys-Adaptive. NUTS shows a pronounced decay with dimension—many models above $d\!\approx\!10$ cluster below $0.05$. In contrast, all three MAKLA variants remain essentially flat across dimension. Coupled MAKLA (middle-left panel) concentrates in the 0.180.180.270.27 band, comparable to the adaptive variants, with a small number of outliers (e.g., indices 5 and 15) where the absence of time-averaged adaptation yields slightly more conservative steps.
Aggregate gains (Table 5).
Relative to NUTS, Coupled MAKLA improves the geometric-mean median ESS/Grad from $1.71\!\times\!10^{-2}$ to $2.43\!\times\!10^{-1}$ , and the minimum from $1.35\!\times\!10^{-2}$ to $2.10\!\times\!10^{-1}$. Adaptive MAKLA variants improve these further to $2.92\!\times\!10^{-1}$ (1sys) and $3.06\!\times\!10^{-1}$ (2sys) for the median, and $2.44\!\times\!10^{-1}$ (1sys) and $2.57\!\times\!10^{-1}$ (2sys) for the minimum. Hence, both coupled and adaptive variants of MAKLA yield order-of-magnitude gains over NUTS. The gains in the geometric mean emphasize that these improvements persist across the entire suit, not just a few easy instances.
[…]
Posterior-mean accuracy.
Figure 3 reports the maximum coordinate-wise absolute relative error in the posterior-mean estimate versus dimension. Across all 45 models, Coupled, 1sys, and 2sys-Adaptive MAKLA achieve errors largely between $0.01$ to $0.03$ under the shared gradient budget, with no systematic degradation as dimension increases. Coupled MAKLA is broadly comparable to the adaptive variants. The maximum relative error of around $0.025$ exhibited by all methods is reasonable given that the sample sizes were $4000\cdot 20=80000$, and this error is expected to scale as $O(1/\sqrt{N})$. Hence no bias can be seen based on these experiments.
Convergence diagnostics.
To ensure efficiency comparisons are not confounded by lack of convergence to the target distribution, we computed parameter-wise Gelman-Rubin diagnostics $\widehat{R}$ for every scalar parameter after warmup and pooled the values across all models (Fig. 4). In our calculation of $\widehat{R}$, for each parameter, we considered samples from each particle (20 particles for adaptive methods, while $8d$ particles for Coupled MAKLA) and repeated each experiment 4 times (so there were 80 parallel chains in total for the adaptive variants, and $32d$ parallel chains for Coupled MAKLA). If some of these chains had chosen a different path, that would have been detected by the $\widehat{R}$ diagnostics. All methods concentrate extremely near $\widehat{R}=1$ (most values <1.01<\!1.01), well within standard thresholds (e.g., $1.05$). Notably, Coupled MAKLA exhibits the tightest concentration around 1, reflecting excellent cross-chain agreement; the adaptive variants show slightly broader, but still narrow, right tails.
Interpretation.
Coupled MAKLA delivers order-of-magnitude gains over NUTS by exploiting spatial averaging, and the adaptive variants achieve similar per-gradient efficiency, often within $10\text{–}25$% of coupled MAKLA’s ESS/Grad, while using far fewer particles ($N=20$ chains versus $8d$ particles for Coupled). Under the two-system lens, this is natural: adaptive MAKLA replaces particle averaging with historical averaging. Instead of computing the preconditioner from a large, simultaneous ensemble, it estimates the same quantity from the chain’s time history. Thus, coupled and adaptive are two ways to approximate the same mean-field preconditioner; one by averaging across particles at a fixed time, the other by averaging across time with many fewer particles. Practically, this trades memory and ensemble size for a lightweight adaptation schedule, retaining most of the efficiency benefits of the coupled ensemble while substantially reducing particle count.
Both designs parallelize trivially: gradients, OABAO updates, and accept/rejects run independently across chains/particles; the only coordination per iteration is a small reduction (one/two covariance estimates) and a broadcast. Short, fixed-length steps ($k=1$) give regular kernels with minimal sync. Coupled MAKLA can saturate the device with its larger batch ($8d$ particles); the adaptive variants use fewer particles ($N=20$), reducing memory/bandwidth while still benefiting from batched linear algebra and vectorized kernels. In both cases, avoiding long serial trajectories makes the methods well-suited to high-throughput GPU/SMID implementations.
Table 5: Geometric means (across all benchmarked posteriors) of median ESS per gradient and minimum ESS per gradient for each sampler (higher is better). Bold indicates the best value in each row.
NUTS Coupled 1sys-Adaptive 2sys-Adaptive
Median ESS/Grad 1.71e-02 2.43e-01 2.92e-01 3.06e-01
Minimum ESS/Grad 1.35e-02 2.10e-01 2.44e-01 2.57e-01
Figure 2: Median ESS/Grad vs. dimension on 45 posteriordb. Each dot is one posterior; indices (1--45) map to Appendix Table 7. The coupled and adaptive MAKLA variants maintain high, nearly flat ESS/Grad across dimensions, whereas NUTS degrades noticeably as dimension increases.
Figure 3: Posterior-mean accuracy vs. dimension on 45 posteriordb. For each of the 45 posteriors, we plot the maximum coordinate-wise absolute relative error (MCARE) in the posterior-mean estimate, $\max_{j}|\hat{\mu}_{j}-\mu_{j}^{\star}|/\text{std}(\mu_{j}^{*})$, against the dimension; the $y$-axis is on a log scale. Reference means $\mu^{\star}$ and standard deviation $\text{std}(\mu_{j}^{*})$ are computed from the gold-standard reference draws distributed with posteriordb; indices (1–45) map to Appendix B Table 7.
Figure 4: Histogram of parameter-wise $\widehat{R}$ (Gelman–Rubin statistics) values across 45 posteriordb. For each of the 45 models and for every scalar parameter, we compute $\hat{R}$ after warmup and pool all values into a single distribution for three samplers: Coupled MAKLA (purple), 1sys-Adaptive MAKLA (green), and 2sys-Adaptive MAKLA (red). All methods concentrate extremely close to the ideal $\widehat{R}=1$ (note the tight axis range), indicating good cross-chain mixing; lower is better.
7.3 Higher dimensional examples
All of the examples with reference samples in posteriordb contain fewer than 100 dimensions. To show the relevance of our method for higher-dimensional sampling problems, we have also tested two additional examples from posteriordb: radon_all-radon_variable_intercept_slope_noncentered (777 dimensions) and three_men3-ldaK2 (505 dimensions). Both of these problems also exhibited poor conditioning and required a longer burn-in period to achieve optimal efficiency. We have implemented the 2sys-Adaptive MAKLA for these problems with 20 particles, using step size $h=0.25$, with 48 00048\,000 burn-in steps followed by 48 00048\,000 sampling steps (thinning was chosen as $4$). As we can see based on the results in Table 6, our sampler exhibits excellent efficiency even for such higher-dimensional problems.
Table 6: Median and minimum ESS/grad for 2sys-Adaptive MAKLA on two high-dimensional models in posteriordb.
radon_all-r_v_i_s_noncent (777 dims) three_men3-ldaK2 (505 dims)
Median ESS/Grad 7.88e-02 3.74e-02
Minimum ESS/Grad 2.70e-02 2.85e-02
Max $\widehat{R}$ (all components) 1.0008 1.0020
8 Conclusion
We introduced a two-system paradigm for constructing finite-particle approximations of mean-field samplers and showed how it yields practical, parallel MCMC algorithms with correct invariance properties. The key idea is to run two interacting ensembles whose proposal statistics are computed cross-system; this symmetric coupling preserves the target law in the mean-field limit and leads, after particle discretization, to a discrete-time ensemble-chain sampler that leaves $\rho^{\otimes 2N}$ invariant while requiring only two covariance evaluations per step.
Our framework clarifies the relationship between three major threads in modern MCMC. Ensemble-chain methods arise as finite approximations of mean-field dynamics; their empirical statistics $\Theta(\delta_{\mathbf{X}_{t}})$ converge to the mean-field quantities $\Theta(\mu_{t})$ as $N\to\infty$. Adaptive samplers recover the same dynamics in the long run by replacing particle averages with time averages. The two-system view presents these within a single framework and motivates concrete algorithms (Coupled MALA and Coupled MAKLA), leveraging cross-system information while preserving the correct stationary law.
In an extensive benchmark of 45 posterior distributions from posteriordb, our adaptive two-system MAKLA variants achieve consistently higher median ESS/grad than NUTS across a wide range of dimensions, indicating dimension-robust per-gradient efficiency. We emphasize that these gains stem from short, SIMD-friendly updates rather than long serial trajectories, making the methods well-suited to single-CPU and GPU execution.
By coupling two interacting systems, we avoid small step size heuristics that plague finite-particle corrections, and we bypass $O(N)$ per-step covariance costs required by self-exclusion schemes, while retaining exact invariance of the desired target. The resulting algorithms inherit the interpretability of mean-field samplers, the efficiency of ensemble methods, and the adaptivity and practicality of single-chain schemes.
We release an open-source Python implementation (pyMALB [^3] https://github.com/pyMALB-Dev/pyMALB and pyMALBSTAN [^4] https://github.com/pyMALB-Dev/pyMALBSTAN) with scripts to reproduce all figures, including the posterior benchmark suite, to facilitate adoption and further comparison.
In sum, divide, interact, and sample provides a principled and scalable route to high-throughput Bayesian computation: it preserves the right target, maps cleanly onto modern hardware, and delivers robust per-gradient efficiency on real-world posteriors.
9 Acknowledgements
We thank Dr. Anna Lisa Varri, Dr. Amanda Lenzi, Prof. Peter Radchenko, Dr. Gonçalo dos Reis, and Dr. Deven Sethi for the helpful conversations while writing this paper. James Chok acknowledges the use of the Isaac Newton Trust (INT) grant G101121 LEAG/929. Daniel Paulin’s research was supported by an NTU Start Up Grant.
Appendix A Appendix A: Detailed Proofs
In the following proofs, we use the notation that $c$ is a constant that changes from line to line and does not depend on the number of particles $N$.
A.1 Drift and Diffusion are square-integrable
Before we begin the proof, we must first state a corollary
Corollary 0.1 (Theorem 1.2 [7]).
Let $\mathbf{X}=(X_{t})_{0\leq t\leq T}$ be a square-integrable process, and define the new process
$U(\mathbf{X})_{t}\ =\ X_{0}\ +\ \int^{t}_{0}b(s,X_{s})\,ds\ +\ \int^{t}_{0}\sigma(s,X_{s})\,dW_{s}.$ (38)
If
$\mathbb{E}\int^{T}_{0}\bigg|\int^{t}_{0}b(s,X_{s})\bigg|^{2}dt\ <\ \infty,\qquad\text{and}\qquad\mathbb{E}\int^{T}_{0}\bigg|\int^{t}_{0}\sigma(s,X_{s})\bigg|^{2}dt\ <\ \infty,$ (39)
Then $U(\mathbf{X})$ is square-integrable and defines a strict contraction. Thus $\mathbf{X}$ is a solution to this SDE.
Theorem 1.
Let $m\in\mathcal{P}_{2}(C([0,T];\mathbb{R}))$ be temporarily fixed, and define the natural bijection $m\mapsto m_{t}$, its time marginals, which is continuous with respect to the Wasserstein distance (Theorem 1.7 [7]). Let $\mathbf{Z}=(Z_{t})_{0\leq t\leq T}$ be a square-integrable process, and define the new process
$U(\mathbf{Z})_{t}\ =\ Z_{0}\ +\ \int^{t}_{0}\mathbf{b}(s,Z_{s},m_{t})\,ds\ +\ \int^{t}_{0}\bm{\sigma}(s,Z_{s},m_{t})\,dW_{s}.$ (40)
Then
$\mathbb{E}\bigg|\int^{t}_{0}\mathbf{b}(s,Z_{s},m_{s})\,ds\bigg|^{2}\,dt\ <\ \infty\qquad\text{and}\qquad\mathbb{E}\bigg|\int^{t}_{0}\bm{\sigma}(s,Z_{s},m_{s})\,ds\bigg|^{2}\,dt\ <\ \infty,$ (41)
and $U(\mathbf{Z})_{t}$ is a strict contraction.
Proof.
By the Lipschitz assumption on $b$
$\lvert b(t,X_{t},\rho_{t})\rvert^{2}\$$\leq\ 2(\lvert b(t,X_{t},\rho_{t})-b(t,0,\rho_{t})\rvert^{2}\ +\ \lvert b(t,0,\rho_{t})\rvert^{2})$ (42)
$\leq c(\lvert X_{t}\rvert^{2}\ +\ \lvert b(t,0,\rho_{t})\rvert^{2}),$ (43)
Thus,
$\lvert\mathbf{b}(t,Z_{t},m_{t})\rvert^{2}\ \leq\ c(\lvert Z_{t}\rvert^{2}+\lvert b(t,0,p_{1}\circ m_{t})\rvert^{2}+\lvert b(t,0,p_{2}\circ m_{t})\rvert^{2}).$ (44)
Using Jensen’s inequality
$\mathbb{E}\bigg|\int^{t}_{0}\mathbf{b}(s,Z_{s},m_{s})\,ds\bigg|^{2}\,dt\$$\leq\ \mathbb{E}\,t\int^{t}_{0}\lvert\mathbf{b}(s,Z_{s},m_{s})\rvert^{2}\,ds\,dt$ (45)
$\leq\ cT\mathbb{E}\int^{t}_{0}(\lvert Z_{s}\rvert^{2}+\lvert b(s,0,p_{1}\circ m_{s})\rvert^{2}+\lvert b(s,0,p_{2}\circ m_{s})\rvert^{2})ds\,dt$ (46)
$\leq\ cT\left(\mathbb{E}\sup_{0\leq t\leq T}\lvert Z_{t}\rvert^{2}\ +\ \mathbb{E}\int^{T}_{0}(\lvert b(s,0,p_{1}\circ m_{s})\rvert^{2}+\lvert b(s,0,p_{2}\circ m_{s})\rvert^{2})ds\right)$ (47)
$<\ \infty,$ (48)
using the fact that $b$ and $Z_{t}$ are square-integrable processes. Similarly, it can be shown that
$\mathbb{E}\bigg|\int^{t}_{0}\bm{\sigma}(s,Z_{s},m_{s})\,ds\bigg|^{2}\,dt\ <\ \infty.$ (49)
Hence, $\mathbf{b}$ and $\bm{\sigma}$ are square-integrable processes. It follows that
$\mathbb{E}\int^{T}_{0}\bigg|\int^{t}_{0}\mathbf{b}(s,Z_{s},m_{s})\bigg|^{2}\ <\ \infty\qquad\text{and}\qquad\mathbb{E}\int^{T}_{0}\bigg|\int^{t}_{0}\bm{\sigma}(s,Z_{s},m_{s})\,ds\bigg|^{2}\,dt\ <\ \infty$ (50)
By Corollary 0.1, $U(\mathbf{Z})_{t}$ is squared-integrable, with $\mathbf{Z}$ the unique solution to this SDE. ∎
A.2 Existence and uniqueness
This proof follows the standard argument of the existence and uniqueness of McKean--Vlasov SDEs, see [7, 21, 42] for more details.
Theorem 2.
For any $T>0$, solutions to the SDEs defined in (18) exist and are unique on $[0,T]$.
Proof.
Using the Lipschitz assumption on $b$,
$\lvert\mathbf{b}(t,Z_{t},\pi_{t})-\mathbf{b}(t,\overline{Z}_{t},\overline{\pi}_{t})\rvert^{2}\$$\leq\ 2c\Big(\lvert Z_{t}-\overline{Z}_{t}\rvert^{2}\ +\ \mathcal{W}_{2}^{2}(p_{1}\circ\pi_{t},p_{1}\circ\overline{\pi}_{t})+\mathcal{W}_{2}^{2}(p_{2}\circ\pi_{t},p_{2}\circ\overline{\pi}_{t})\Big)$ (51)
Similarly,
$\lvert\bm{\sigma}(t,Z_{t},\pi_{t})-\bm{\sigma}(t,\overline{Z}_{t},\overline{\pi}_{t})\rvert\ \leq\ 2c\Big(\lvert Z_{t}-\overline{Z}_{t}\rvert^{2}\ +\ \mathcal{W}_{2}^{2}(p_{1}\circ\pi_{t},p_{1}\circ\overline{\pi}_{t})+\mathcal{W}_{2}^{2}(p_{2}\circ\pi_{t},p_{2}\circ\overline{\pi}_{t})\Big).$ (52)
For $t\in[0,T]$, Jensen’s inequality yields
$\lvert Z_{s}-\overline{Z}_{s}\rvert^{2}\ \leq\ 2t\int^{t}_{0}\lvert\mathbf{b}(Z_{s},\pi_{s})-\mathbf{b}(\overline{Z}_{s},\overline{\pi}_{s})\rvert^{2}\,dr\ +\ 2\bigg|\int^{t}_{0}(\bm{\sigma}(Z_{s},\pi_{s})-\bm{\sigma}(\overline{Z}_{s},\overline{\pi}_{s})dW_{s}\bigg|^{2}$ (53)
Thus, using Doob’s maximal inequality followed by Itô’s isometry, and the Lipschitz assumption
$\mathbb{E}\sup_{0\leq s\leq t}\lvert Z_{s}-\overline{Z}_{s}\rvert^{2}\$$\leq\ 2T\mathbb{E}\sup_{0\leq s\leq t}\int_{0}^{s}\lvert\mathbf{b}(r,Z_{r},\pi_{r})-\mathbf{b}(r,\overline{Z}_{r},\overline{\pi}_{r})\rvert^{2}dr$ (54)
+ $\qquad\qquad+\ 2\mathbb{E}\sup_{0\leq s\leq t}\bigg|\int^{s}_{0}[\bm{\sigma}(r,Z_{r},\pi_{r})-\bm{\sigma}(r,\overline{Z}_{r},\overline{\pi}_{r})]dW_{r}\bigg|^{2}$ (55)
$\leq\ 2T\mathbb{E}\int_{0}^{t}\lvert\mathbf{b}(r,Z_{r},\pi_{r})-\mathbf{b}(r,\overline{Z}_{r},\overline{\pi}_{r})\rvert^{2}dr$ (56)
+ $\qquad\qquad+\ 8\mathbb{E}\int^{t}_{0}\lvert\bm{\sigma}(r,Z_{r},\pi_{r})-\bm{\sigma}(r,\overline{Z}_{r},\overline{\pi}_{r})\rvert^{2}dr$ (57)
$\leq c\bigg(\int^{t}_{0}\mathbb{E}\sup_{0\leq r\leq s}\lvert Z_{r}-\overline{Z}_{r}\rvert^{2}ds$ (58)
$\qquad\qquad+\int^{t}_{0}\mathcal{W}_{2}^{2}(p_{1}\circ\pi_{s},p_{1}\circ\overline{\pi}_{s})+\mathcal{W}_{2}^{2}(p_{2}\circ\pi_{s},p_{2}\circ\overline{\pi}_{s})\,ds\bigg).$ (59)
By Theorem 1, $(Z_{t})_{0\leq t\leq T}$ is a square-integrable process, and therefore bounded. Thus, Grönwalls inequality yields
$\mathbb{E}\sup_{0\leq s\leq t}\lvert Z_{t}-\overline{Z}_{t}\rvert^{2}\ \leq\ ce^{cT}\int^{t}_{0}\mathcal{W}_{2}^{2}(p_{1}\circ\pi_{s},p_{1}\circ\overline{\pi}_{s})+\mathcal{W}_{2}^{2}(p_{2}\circ\pi_{s},p_{2}\circ\overline{\pi}_{s})\,ds.$ (60)
This lets us perform the standard contraction argument to prove uniqueness and existence. For completeness, we provide it here.
Define $\Phi$ the map which associates $m\in\mathcal{P}_{2}(C([0,T];\mathbb{R}^{2d})$, the space of continuous measures on $\mathbb{R}^{2d}$, to the law of the solution of
$Z_{t}\ =\ Z_{0}\ +\ \int^{t}_{0}\mathbf{b}(s,Z_{s},m_{s})\,ds\ +\ \int^{t}_{0}\bm{\sigma}(s,Z_{s},m_{s})\,dW_{s},$ (61)
where $m\mapsto m_{s}$ is the continuous natural bijection map with respect to the Wasserstein distance.
Then a process $(Z_{t})_{0\leq t\leq T}$ satisfying $\mathbb{E}\sup_{0\leq t\leq T}\lvert Z_{t}\rvert^{2}<\infty$ is a solution of (18) if and only if its law is a fixed point of $\Phi$. Define the Wasserstein distance on $\mathcal{P}_{2}(C([0,T];\mathbb{R}^{2d})$ as
$\mathcal{W}_{t,2}(m_{1},m_{2})\ =\$$\inf\bigg\{\left(\int^{t}_{0}\sup_{0\leq s\leq t}\lvert X_{s}(w_{1})-X_{s}(w_{2})\rvert^{2}\,m(dw_{1},dw_{2})\right)^{1/2};$ (62)
$\qquad\qquad m\in\mathcal{P}_{2}(C([0,T],\mathbb{R}^{d})\times C([0,T],\mathbb{R}^{d}))\ \text{with marginals}\ m_{1}\text{ and }m_{2}\bigg\}.$
Since the projections $p_{i}\circ\Phi(m)$ is a marginal of $\Phi(m)$, it follows that $p_{i}\circ\Phi(m)=\textrm{Law}((Z_{i,t})_{0\leq t\leq T})$. Hence,
$\sum_{i=1,2}\mathcal{W}^{2}_{t,2}(p_{i}\circ\Phi(m),p_{i}\circ\Phi(\overline{m}))\$$\leq\ \ \mathbb{E}\sup_{0\leq s\leq t}\lvert Z_{1,t}-\overline{Z}_{1,t}\rvert^{2}+\lvert Z_{2,t}-\overline{Z}_{2,t}\rvert^{2},$ (63)
$=\ \mathbb{E}\sup_{0\leq s\leq t}\lvert Z_{t}-\overline{Z}_{t}\rvert^{2}.$ (64)
Noting that $\mathcal{W}_{2}(p_{i}\circ m_{s},p_{i}\circ\overline{m}_{s})\leq\mathcal{W}_{s,2}(p_{i}\circ m,p_{i}\circ\overline{m})$ yields
$\sum_{i=1,2}\mathcal{W}^{2}_{t,2}(p_{i}\circ\Phi(m),p_{i}\circ\Phi(\overline{m}))\ \leq\ ce^{cT}\int^{t}_{0}\Big(\sum_{i=1,2}\mathcal{W}_{s,2}^{2}(p_{i}\circ m,p_{i}\circ\overline{m})\Big)\,ds.$ (65)
Iterating this inequality gives
$\sum_{i=1,2}\mathcal{W}^{2}_{T,2}(p_{i}\circ\Phi^{k}(m),p_{i}\circ\Phi^{k}(\overline{m}))\$$\leq\ c^{k}e^{kcT}\int^{T}_{0}\frac{(T-s)^{k-1}}{(k-1)!}\Big(\sum_{i=1,2}\mathcal{W}_{s,2}^{2}(p_{i}\circ m,p_{i}\circ\overline{m})\Big)\,ds$ (66)
$\leq\ \frac{c^{k}T^{k}}{k!}\sum_{i=1,2}\mathcal{W}_{T,2}^{2}(p_{i}\circ m,p_{i}\circ\overline{m}).$ (67)
Note that a coupling between the marginals of a distribution gives rise to a coupling on the full distribution. Thus,
$\mathcal{W}_{T,2}^{2}(\Phi^{k}(m),\Phi^{k}(\overline{m}))\ \leq\ \mathcal{W}_{T,2}^{2}(p_{i}\circ\Phi^{k}(m),p_{i}\circ\Phi^{k}(\overline{m})).$ (68)
Thus
$\mathcal{W}_{T,2}^{2}(\Phi^{k}(m),\Phi^{k}(\overline{m}))\$$\leq\ \frac{1}{2}\sum_{i=1,2}\mathcal{W}^{2}_{t,2}(p_{i}\circ\Phi^{k}(m),p_{i}\circ\Phi^{k}(\overline{m}))\ \leq\ \frac{c^{k}T^{k}}{k!}\sum_{i=1,2}\mathcal{W}_{T,2}^{2}(p_{i}\circ m,p_{i}\circ\overline{m}),$ (69)
which shows that for $k$ large enough, $\Phi^{k}$ is a strict contraction, and hence $\Phi$ admits $\textrm{Law}((Z_{t})_{0\leq t\leq T})$ as a unique fixed point. ∎
A.3 Convergence in Wasserstein
Corollary 2.1.
If $\textrm{Law}(X_{0})=\textrm{Law}(Y_{0})=\textrm{Law}(\overline{X}_{0})=\textrm{Law}(\overline{Y}_{0})$, then $\textrm{Law}(X_{t})=\textrm{Law}(Y_{t})=\textrm{Law}(\overline{X}_{t})=\textrm{Law}(\overline{Y}_{t})$ for $t\in[0,T]$. That is, the mean-field dynamics of the two-system approach are the same as using one particle.
Proof.
Using Theorem 3, the result follows as the Wasserstein distance defines a metric on probability measures on $\mathbb{R}^{d}$. ∎
Theorem 3.
Consider the coupled systems of SDEs
$\begin{split}&dX_{t}\ =\ b(t,X_{t},\rho_{t})\,dt\ +\ \sigma(t,X_{t},\rho_{t})\,dW_{t}^{X},\\ &dY_{t}\ =\ b(t,Y_{t},\mu_{t})\,dt\ +\ \sigma(t,Y_{t},\mu_{t})\,dW_{t}^{Y},\\ &d\overline{X}_{t}\ =\ b(t,\overline{X}_{t},\overline{\mu}_{t})\,dt\ +\ \sigma(t,\overline{X}_{t},\overline{\mu}_{t})\,dW_{t}^{X}\\ &d\overline{Y}_{t}\ =\ b(t,\overline{Y}_{t},\overline{\rho}_{t})\,dt\ +\ \sigma(t,\overline{Y}_{t},\overline{\rho}_{t})\,dW_{t}^{Y}\end{split}$ (70)
with $\textrm{Law}(X_{t})=\mu_{t},\,\textrm{Law}(Y_{t})=\rho_{t},\,\textrm{Law}(\overline{X}_{t})=\overline{\mu}_{t},\,\textrm{Law}(\overline{Y}_{t})=\overline{\rho}_{t}$ and initial conditions $X_{0},Y_{0},\overline{X}_{0},\overline{Y}_{0}$ respectively. Then for $t\in[0,T]$,
$\mathcal{W}_{2}^{2}(\mu_{t},\overline{\mu}_{t})\ +\ \mathcal{W}_{2}^{2}(\rho_{t},\overline{\mu}_{t})\ \leq\ c\Big[\mathcal{W}_{2}^{2}(\mu_{0},\overline{\mu}_{0})\ +\ \mathcal{W}_{2}^{2}(\rho_{0},\overline{\rho}_{0})\ +\ \mathcal{W}_{2}^{2}(\overline{\mu}_{0},\overline{\rho}_{0})\Big].$ (71)
where $c$ is a constant that depends on $t$, and $\mathcal{W}_{2}$ is the Wasserstein distance. In other words, if the two system start close (in Wasserstein distance), then they remain close for $t\in[0,T]$.
Proof.
For $t\in[0,T]$, Itô’s isometry, Jensen’s inequality, and the Lipschitz assumption,
$\mathbb{E}\lvert X_{t}-\overline{X}_{t}\rvert^{2}\$$\leq\ c\mathbb{E}\lvert X_{0}-\overline{X}_{0}\rvert\ +\ c\mathbb{E}\bigg|\int^{t}_{0}(b(s,X_{s},\rho_{s})-b(s,\overline{X}_{s},\overline{\mu}_{s}))\,ds\bigg|^{2}$ (72)
$\qquad\qquad+\ c\mathbb{E}\int^{t}_{0}\lvert\sigma(s,X_{s},\rho_{s})-\sigma(s,\overline{X}_{s},\overline{\mu}_{s})\rvert^{2}\,ds$ (73)
$\leq\ c\mathbb{E}\lvert X_{0}-\overline{X}_{0}\rvert^{2}\ +\ c\int^{t}_{0}\mathbb{E}\lvert b(s,X_{s},\rho_{s})-b(s,\overline{X}_{s},\overline{\mu}_{s})\rvert^{2}\,ds$ (74)
$\qquad\qquad+\ c\int^{t}_{0}\mathbb{E}\lvert\sigma(s,X_{s},\rho_{s})-\sigma(s,\overline{X}_{s},\overline{\mu}_{s})\rvert^{2}\,ds$ (75)
$\leq\ c\mathbb{E}\lvert X_{0}-\overline{X}_{0}\rvert^{2}\ +\ c\int^{t}_{0}\Big(\mathbb{E}\lvert X_{s}-\overline{X}_{s}\rvert^{2}\ +\ \mathcal{W}_{2}^{2}(\rho_{s},\overline{\mu}_{s})\Big)\,ds.$ (76)
By Theorem 1, $(X_{t})_{0\leq t\leq T}$, and $(\overline{X}_{t})_{0\leq t\leq T}$ are squared-integrable, it follows that $\lvert X_{t}-\overline{X}_{t}\rvert$ is also squared-integrable. Thus, by Grönwall’s lemma
$\mathbb{E}\lvert X_{t}-\overline{X}_{t}\rvert^{2}\$$\leq\ e^{cT}\mathbb{E}\lvert X_{0}-\overline{X}_{0}\rvert^{2}\ +\ e^{cT}\int^{t}_{0}\mathcal{W}_{2}^{2}(\rho_{s},\overline{\mu}_{s})\,ds.$ (77)
Since this holds for any couplings of $X_{t}$ and $\overline{X}_{t}$, taking the infimum of the left side yields
$\mathcal{W}_{2}^{2}(\mu_{t},\overline{\mu}_{t})\ \leq\ e^{cT}\mathbb{E}\lvert X_{0}-\overline{X}_{0}\rvert^{2}\ +\ e^{cT}\int^{t}_{0}\mathcal{W}_{2}^{2}(\rho_{s},\overline{\mu}_{s})\,ds.$ (78)
Since a similar inequality holds for $\mathbb{E}\lvert Y_{t}-\overline{Y}_{t}\rvert^{2}$,
$\mathcal{W}_{2}^{2}(\mu_{t},\overline{\mu}_{t})\ +\ \mathcal{W}_{2}^{2}(\rho_{t},\overline{\rho}_{t})\$$\leq\ c\left(\mathbb{E}\lvert X_{0}-\overline{X}_{0}\rvert^{2}\ +\ \mathbb{E}\lvert Y_{0}-\overline{Y}_{0}\rvert^{2}\right)$ (79)
$\qquad\qquad+c\int^{t}_{0}\Big(\mathcal{W}_{2}^{2}(\rho_{s},\overline{\mu}_{s})\ +\ \mathcal{W}_{2}^{2}(\mu_{s},\overline{\rho}_{s})\Big)ds.$ (80)
$\leq\ c\left(\mathbb{E}\lvert X_{0}-\overline{X}_{0}\rvert^{2}\ +\ \mathbb{E}\lvert Y_{0}-\overline{Y}_{0}\rvert^{2}\right)$ (81)
$\qquad\qquad+c\int^{t}_{0}\Big(\mathcal{W}_{2}^{2}(\rho_{s},\overline{\rho}_{s})\ +\ \mathcal{W}_{2}^{2}(\mu_{s},\overline{\mu}_{s})\ +\ \mathcal{W}_{2}^{2}(\overline{\rho}_{s},\overline{\mu}_{s})\Big)ds,$ (82)
using the triangle inequality. Since $(X_{t})_{0\leq t\leq T}$, $(Y_{t})_{0\leq t\leq T}$, $(\overline{X}_{t})_{0\leq t\leq T}$, and $(\overline{Y}_{t})_{0\leq t\leq T}$ are squared-integrable, it follows that $\mathbb{E}\lvert X_{t}-\overline{X}_{t}\rvert^{2}$ and $\mathbb{E}\lvert Y_{t}-\overline{Y}_{t}\rvert^{2}$ are bounded for $t\in[0,T]$. Hence $\mathcal{W}_{2}^{2}(\mu_{t},\overline{\mu}_{t})+\mathcal{W}_{2}^{2}(\rho_{t},\overline{\rho}_{t})+\mathcal{W}_{2}^{2}(\overline{\mu}_{t},\overline{\rho}_{t})$ is also bounded on $t\in[0,T]$. This allows us to apply Grönwall’s inequality
$\mathcal{W}_{2}^{2}(\mu_{t},\overline{\mu}_{t})\ +\ \mathcal{W}_{2}^{2}(\rho_{t},\overline{\rho}_{t})\ \leq\ c\left(\mathbb{E}\lvert X_{0}-\overline{X}_{0}\rvert^{2}\ +\ \mathbb{E}\lvert Y_{0}-\overline{X}_{0}\rvert^{2}\ +\ \int^{t}_{0}\mathcal{W}^{2}_{2}(\overline{\rho}_{s},\overline{\mu}_{s})\,ds\right).$ (83)
A similar argument can be used to show that $\mathcal{W}^{2}_{2}(\overline{\mu}_{t},\overline{\rho}_{t})\leq c\mathbb{E}|\overline{X}_{0}-\overline{Y}_{0}|^{2}$. Thus,
$\mathcal{W}_{2}^{2}(\mu_{t},\overline{\mu}_{t})\ +\ \mathcal{W}_{2}^{2}(\rho_{t},\overline{\rho}_{t})\ \leq\ c\left(\mathbb{E}\lvert X_{0}-\overline{X}_{0}\rvert^{2}\ +\ \mathbb{E}\lvert Y_{0}-\overline{X}_{0}\rvert^{2}\ +\ \mathbb{E}\lvert\overline{Y}_{0}-\overline{X}_{0}\rvert^{2}\right).$ (84)
Since this holds for any couplings of the initial condition, taking the infimum of the right side of this inequality yields
$\mathcal{W}_{2}^{2}(\mu_{t},\overline{\mu}_{t})\ +\ \mathcal{W}_{2}^{2}(\rho_{t},\overline{\mu}_{t})\ \leq\ c\Big[\mathcal{W}_{2}^{2}(\mu_{0},\overline{\mu}_{0})\ +\ \mathcal{W}_{2}^{2}(\rho_{0},\overline{\rho}_{0})\ +\ \mathcal{W}_{2}^{2}(\overline{\mu}_{0},\overline{\rho}_{0})\Big].$ (85)
A.4 Propagation of Chaos
Before proving the propagation of chaos, we first state an important Lemma which measures how ‘far’ independent and identically distributed samples from a distribution, $\mu$, are away from $\mu$.
Lemma 2 (Lemma 1.9 [7]).
Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $\bm{\xi}_{N}=(\xi_{1},\ldots\xi_{N})$ be a sequence of independent random variables with common law $\mu$. Then for each $N\geq 1$, we have
$\mathbb{E}\mathcal{W}_{2}^{2}(\delta_{\bm{\xi}_{N}},\mu)\ \leq\ 4\int_{\mathbb{R}^{d}}\lvert{x}\rvert^{2}\,\mu(dx),\qquad\text{and}\qquad\lim_{N\to\infty}\mathbb{E}\mathcal{W}_{2}^{2}(\delta_{\bm{\xi}_{N}},\mu)\ =\ 0.$ (86)
Theorem 4 (Propagation of Chaos).
Under assumptions (1) and (2), the coupled $2N$-particle system (20) approaches the mean-field limit as $N\to\infty$ in the following sense: for the coupled SDEs,
$\begin{split}dX^{i}_{t}\ &=\ b(t,X^{i}_{t},\delta_{\mathbf{Y}_{t}})\,dt\ +\ \sigma(t,X^{i}_{t},\delta_{\mathbf{Y}_{t}})\,dW^{X,i}_{t}\\ dY^{i}_{t}\ &=\ b(t,Y^{j}_{t},\delta_{\mathbf{X}_{t}})\,dt\ +\ \sigma(t,Y^{i}_{t},\delta_{\mathbf{X}_{t}})\,dW^{Y,i}_{t}\\ d\overline{X}^{i}_{t}\ &=\ b(t,\overline{X}^{i}_{t},\overline{\mu}_{t})\,dt\ +\ \sigma(t,\overline{X}^{i}_{t},\overline{\mu}_{t})\,dW^{X,i}_{t},\\ d\overline{Y}^{i}_{t}\ &=\ b(t,\overline{Y}^{i}_{t},\overline{\rho}_{t})\,dt\ +\ \sigma(t,\overline{Y}^{i}_{t},\overline{\rho}_{t})\,dW^{Y,i}_{t},\end{split}$ (87)
for $i\in\{1,\ldots,N\}$, with identical initial conditions $X^{i}_{0}=Y^{i}_{0}=\overline{X}^{i}_{0}=\overline{Y}^{i}_{0}=X_{0}$, we have
$\lim_{N\to\infty}\sup_{1\leq i\leq N}\mathbb{E}\left(\sup_{0\leq t\leq T}\lvert X^{i}_{s}-\overline{X}^{i}_{s}\rvert^{2}+\sup_{0\leq t\leq T}\lvert Y^{i}_{s}-\overline{Y}^{i}_{s}\rvert^{2}\right)\ =\ 0,$ (88)
with a convergence rate
$\sup_{1\leq i\leq N}\left(\mathbb{E}\sup_{0\leq t\leq T}\lvert X^{i}_{t}-\overline{X}^{i}_{t}\rvert^{2}\ +\ \mathbb{E}\sup_{0\leq t\leq T}\lvert Y^{i}_{t}-\overline{Y}^{i}_{t}\rvert^{2}\right)\ \leq\ \frac{c}{N},$ (89)
where $c$ is a constant that does not depend on $N$.
Proof.
This proof follows Theorem 1.10 from [7]. By Doob’s maximal inequality and Itô’s isometry
$\mathbb{E}\sup_{0\leq s\leq t}\lvert X^{i}_{s}-\overline{X}^{i}_{s}\rvert^{2}\$$\leq\ c\mathbb{E}\bigg|\int^{t}_{0}(b(s,X^{i}_{s},\delta_{\mathbf{Y}_{s}})-b(s,\overline{X}^{i}_{s},\overline{\mu}_{s}))ds\bigg|^{2}$ (90)
$\qquad\qquad+\ c\mathbb{E}\int^{t}_{0}\lvert\sigma(s,X^{i}_{s},\delta_{{\mathbf{Y}}_{s}})-\sigma(s,\overline{X}^{i}_{s},\overline{\mu}_{s})\rvert^{2}\,ds$ (91)
$\begin{split}&\leq\ c\int^{t}_{0}\mathbb{E}\lvert b(s,X^{i}_{s},\delta_{\mathbf{Y}_{s}})-b(s,\overline{X}^{i}_{s},\delta_{\overline{\mathbf{X}}_{s}})\rvert^{2}\,ds\\ &\qquad\qquad+\ c\int^{t}_{0}\mathbb{E}\lvert b(s,\overline{X}^{i}_{s},\delta_{\overline{\mathbf{X}}_{s}})-b(s,\overline{X}^{i}_{s},\overline{\mu}_{s})\rvert^{2}\,ds\\ &+\ c\int^{t}_{0}\mathbb{E}\lvert\sigma(s,X^{i}_{s},\delta_{\mathbf{Y}_{s}})-\sigma(s,\overline{X}^{i}_{s},\delta_{\overline{\mathbf{X}}_{s}})\rvert^{2}\,ds\\ &\qquad\qquad+\ c\int^{t}_{0}\mathbb{E}\lvert\sigma(s,\overline{X}^{i}_{s},\delta_{\overline{\mathbf{X}}_{s}})-\sigma(s,\overline{X}^{i}_{s},\overline{\mu}_{s})\rvert^{2}\,ds,\end{split}$ (92)
with an similar inequality holding for $\mathbb{E}\sup_{0\leq s\leq t}\lvert Y^{j}_{s}-\overline{Y}^{j}_{s}\rvert^{2}$.
Using the triangle inequality, and using a bound on the Wasserstein distance between two Dirac measures (Equation 1.24 [7])
$\mathcal{W}_{2}^{2}(\delta_{\mathbf{Y}_{s}},\delta_{\overline{\mathbf{X}}_{s}})\ \leq\ \mathcal{W}_{2}^{2}(\delta_{\mathbf{Y}_{s}},\delta_{\overline{\mathbf{Y}}_{s}})+\mathcal{W}_{2}^{2}(\delta_{\overline{\mathbf{Y}}_{s}},\delta_{\overline{\mathbf{X}}_{s}})\leq\ \frac{1}{N}\sum_{i=1}^{N}\lvert Y_{s}^{i}-\overline{Y}_{s}^{i}\rvert^{2}+\mathcal{W}_{2}^{2}(\delta_{\overline{\mathbf{Y}}_{s}},\delta_{\overline{\mathbf{X}}_{s}}).$ (93)
Since the random processes $\overline{\mathbf{Y}}_{t}$ and $\overline{\mathbf{X}}_{t}$ start with the same initial condition, existence and uniqueness of strong solutions to the McKean–Vlasov process implies uniquness in law, $\mathcal{W}_{2}^{2}(\delta_{\overline{\mathbf{Y}}_{s}},\delta_{\overline{\mathbf{X}}_{s}})=0$. Thus, by the exchangeability of the couples $(Y^{i},\overline{Y}^{i})$, the Lipschitz assumption of $b$ and $\sigma$ bounds the first and third term of (LABEL:appendix:reandom_eqquation) by
$c\int^{t}_{0}\left(\mathbb{E}\sup_{0\leq r\leq s}\lvert X_{r}^{i}-\overline{X}^{i}_{r}\rvert^{2}\ +\ \mathbb{E}\sup_{0\leq r\leq s}\lvert Y^{i}_{r}-\overline{Y}^{i}_{r}\rvert^{2}\right)\,ds.$ (94)
Combining the inequality for $\mathbb{E}\sup_{0\leq s\leq t}\lvert Y^{j}_{s}-\overline{Y}^{j}_{s}\rvert^{2}$, Grönwall’s lemma yields and the Lipschitz assumption yields
$\mathbb{E}\sup_{0\leq t\leq T}\lvert X^{i}_{t}-\overline{X}^{i}_{t}\rvert^{2}\ +\ \mathbb{E}\sup_{0\leq t\leq T}\lvert Y^{j}_{t}-\overline{Y}^{j}_{t}\rvert^{2}\$$\leq\ c\int^{T}_{0}\mathbb{E}\lvert b(s,\overline{X}^{i}_{s},\delta_{\overline{\mathbf{X}}_{s}})-b(s,\overline{X}^{i}_{s},\overline{\mu}_{s})\rvert^{2}\,ds$ (95)
$\qquad\qquad+\ c\int^{T}_{0}\mathbb{E}\lvert\sigma(s,\overline{X}^{i}_{s},\delta_{\overline{\mathbf{X}}_{s}})-\sigma(s,\overline{X}^{i}_{s},\overline{\mu}_{s})\rvert^{2}\,ds$ (96)
$\quad+\ c\int^{T}_{0}\mathbb{E}\lvert b(s,\overline{Y}^{i}_{s},\delta_{\overline{\mathbf{Y}}_{s}})-b(s,\overline{Y}^{i}_{s},\overline{\rho}_{s})\rvert^{2}\,ds$ (97)
$\qquad\qquad+\ c\int^{T}_{0}\mathbb{E}\lvert\sigma(s,\overline{Y}^{i}_{s},\delta_{\overline{\mathbf{Y}}_{s}})-\sigma(s,\overline{Y}^{i}_{s},\overline{\rho}_{s})\rvert^{2}\,ds$ (98)
$\leq c\int^{T}_{0}\left(\mathbb{E}\mathcal{W}_{2}^{2}(\delta_{\overline{\mathbf{X}}_{s}},\overline{\mu}_{s})\ +\ \mathbb{E}\mathcal{W}_{2}^{2}(\delta_{\overline{\mathbf{Y}}_{s}},\overline{\rho}_{s})\right)\,ds.$ (99)
Since $\overline{X}_{s}^{i},\overline{Y}_{s}^{j}$ are all independent random variables with common law $\overline{\mu}_{s},\overline{\rho}_{s}$, it follows by Lemma 2 and Lebesgue’s dominated convergence theorem that
$\lim_{N\to\infty}\sup_{1\leq i\leq N}\mathbb{E}\left(\sup_{0\leq t\leq T}\lvert X^{i}_{t}-\overline{X}^{i}_{t}\rvert^{2}+\sup_{0\leq t\leq T}\lvert Y^{i}_{t}-\overline{Y}^{i}_{t}\rvert^{2}\right)\ =\ 0.$ (100)
Since $b$ is a linear functional of the measure
$\mathbb{E}\lvert b(s,\overline{X}^{i}_{s},\delta_{\mathbf{Y}_{s}})-b(s,\overline{X}^{i}_{s},\overline{\mu}_{s})\rvert^{2}\$$=\ \mathbb{E}\bigg|b(s,\overline{X}^{i}_{s},\overline{\mu}_{s})-\frac{1}{N}\sum_{j=1}^{N}b(s,\overline{X}^{i}_{s},\delta_{\overline{X}^{j}_{s}})\bigg|^{2}$ (101)
$=\ \frac{1}{N^{2}}\sum_{k,l=1}^{N}\mathbb{E}\bigg(\bigg[\int b(s,\overline{X}^{i}_{s},y)\overline{\mu}_{s}(dy)-b(s,\overline{X}^{i},\delta_{\overline{X}^{k}_{s}})\bigg]$ (102)
$\qquad\qquad\qquad\qquad\cdot\bigg[\int b(s,\overline{X}^{i}_{s},y)\overline{\mu}_{s}(dy)-b(s,\overline{X}^{i},\delta_{\overline{X}^{i}_{s}})\bigg]\bigg).$ (103)
Since $\overline{\mathbf{X}}_{s}$ are all independent random variables with common law $\overline{\mu}_{s}$, the expectation in the summand vanishes when $k\neq l$. Since $b$ is square-integrable, the above equation is bounded by $c/N$ for some constant $c$ that does not depend on $N$. A similar argument follows to bound $\sigma$, thus yielding
$\sup_{1\leq i\leq N}\left(\mathbb{E}\sup_{0\leq t\leq T}\lvert X^{i}_{t}-\overline{X}^{i}_{t}\rvert^{2}\ +\ \mathbb{E}\sup_{0\leq t\leq T}\lvert Y^{i}_{t}-\overline{Y}^{i}_{t}\rvert^{2}\right)\ \leq\ \frac{c}{N}.$ (104)
A.5 Adaptive MALA
We present our algorithm using a single chain and demonstrate that it is geometrically ergodic. The extension to an ensemble of adaptive kernels follows.
The one-chain adaptive MALA considers the proposals
$\widetilde{X}_{n}\ =\ X_{n}\ +\ h\,(C_{n}+\varepsilon I_{d})\,D(X_{n})\ +\ \sqrt{2h(C_{n}+\varepsilon I_{d})}\,\Delta W_{n},$ (105)
equivalently written as
$Q_{C_{n}}(\widetilde{X}_{n}\mid X_{n})\ \propto\ \exp\left(-\frac{1}{4h}\bigg\lVert\left(C_{n}^{-1/2}\ +\ \varepsilon I_{d}\right)\left(\widetilde{X}_{n}\ -\ D(X_{n})\right)\bigg\rVert^{2}\right),$ (106)
for some small $\varepsilon>0$, where, for some $\delta>0$, we bound the drift as
[…]
The adaptive covariance matrix $C_{n}$ is iteratively updated (and bounded) by
$\widetilde{C}_{n}\ =\ \left(1-\frac{a}{n}\right)C_{n-1}\ +\ \frac{a}{n}C(X_{n-1}),\qquad\text{and}\qquad C_{n}\ =\ \frac{K}{\max(K,\lvert\widetilde{C}_{n}\lvert)}\widetilde{C}_{n}.$ (108)
The Metropolis--Hastings acceptance ratio is chosen as
$A_{C_{n}}(X_{n},\widetilde{X}_{n})\ =\ \min\left(1,\ \frac{\rho(\widetilde{X}_{n})Q_{C_{n}}(X_{n}\lvert\widetilde{X}_{n})}{\rho(X_{n})Q_{C_{n}}(\widetilde{X}_{n}\rvert X_{n})}\right).$ (109)
To show geometric ergodicity of this chain, we state the following Corollary.
Corollary 4.1 (Atchadé [1] Proposition 2.1).
Assume the target density $\rho(x)$ has a finite second moment and satisfies the following conditions
$\lim_{\lvert{x}\rvert\to\infty}\frac{x}{\lvert{x}\rvert}\cdot\nabla\log\{\rho(x)\}\ =\ -\infty,\qquad\text{and}\qquad\limsup_{\lvert{x}\rvert\to\infty}\frac{x}{\lvert{x}\rvert}\cdot\frac{\nabla\log\{\rho(x)\}}{\lvert\nabla\log\{\rho(x)\}\rvert}\ <\ 0.$ (110)
For $0<\alpha<1$ write $V_{\alpha}(x)=c_{\alpha}\rho^{-\alpha}(x)$ where $c_{\alpha}$ is such that $V_{\alpha}\geq 1$. Assume that there exists constants $0<\varepsilon_{1}<\varepsilon_{2}<\infty$ and $0<k_{1}<k_{2}<\infty$ such that the proposal kernel $P_{\gamma}(x,y)$ satisfies
$k_{1}g_{\varepsilon_{1}}(y-x)\ \leq\ P_{\gamma}(x,y)\ \leq\ k_{2}g_{\varepsilon_{2}}(y-x),\quad\text{for all}\quad\gamma\in\Gamma,$ (111)
with $\Gamma$ compact, and, for $0<a<\infty$, $g_{a}$ is the density of a $d$-dimensional normal distribution with mean 0 and covariance $aI_{d}$. Then $P_{\gamma}$ is geometrically ergodic.
The proof follows by checking that the proposal kernel (106) satisfies the assumptions in Corollary 4.1.
Theorem 5.
Adaptive MALA defined through Eqns. (106109) is geometrically ergodic and satisfies diminishing adaptation.
Proof.
Since $D(x)$ is bounded, with $\Gamma$ compact and convex, it follows that we can find $k_{1},k_{2}$ and $\varepsilon_{1},\varepsilon_{2}$ such that (111) is satisfied. Moreover, the mapping $C\mapsto P_{C}(x,\cdot)$ is Lipschitz with a diminishing adaptation and therefore is ergodic. ∎
Remark 4.
As noted in [1, 16], for polynomials $h(x),p(x)$ with $h(x)\geq 0$, the assumption on $\log\{\rho(x)\}$ in Corollary 4.1 is satisfied by densities of the form $h(x)e^{-p(x)}$ and $h(x)^{-p(x)}$.
Remark 5.
Atachadé [1] uses the alternate update scheme
$\widetilde{X}^{j}_{n}\ =\ X^{j}_{n}\ +\ hD(X^{j}_{n})\ +\ \sqrt{2h(C_{n}+\varepsilon I_{d})}\Delta W_{n}.$ (112)
That is, the drift is not multiplied by the covariance matrix.
Remark 6.
In [29], Marshall and Roberts show under different assumptions that the tamed drift is unnecessary for the adaptive chain to converge to $\rho$. However, we choose the tamed drift for numerical and theoretical reasons [39, 5, 4].
A.6 Ensemble chain MCMC preserves the correct invariant measure
Theorem 6 (Block MH with a frozen subset preserves $\rho^{\otimes N}$).
Let $\rho$ be a probability density on $\mathbb{R}^{d}$ and let $\pi:=\rho^{\otimes N}$ be the product target on $(\mathbb{R}^{d})^{N}$. Fix a nonempty subset $S\subset\{1,\ldots,N\}$ and write $S^{c}=\{1,\ldots,N\}\setminus S$. For a current state $x=(x_{S},x_{S^{c}})\in(\mathbb{R}^{d})^{N}$, consider a block proposal that keeps the coordinates in $S$ fixed and proposes new values for the coordinates in $S^{c}$ with density
$q_{S}(y_{S^{c}}\mid x_{S^{c}};\,x_{S}),\qquad y_{S^{c}}\in(\mathbb{R}^{d})^{|S^{c}|},$
which may depend arbitrarily on the frozen block $x_{S}$. Define the joint proposal kernel on $(\mathbb{R}^{d})^{N}$ by
$Q_{S}(x,dy)\;=\;\delta_{x_{S}}(dy_{S})\;q_{S}(y_{S^{c}}\mid x_{S^{c}};\,x_{S})\,dy_{S^{c}}.$
Accept the proposed move $x\to y=(x_{S},y_{S^{c}})$ with MH probability
$\alpha_{S}(x,y)\;=\;\min\!\left\{1,\;\frac{\pi(y)\,q_{S}(x_{S^{c}}\mid y_{S^{c}};\,x_{S})}{\pi(x)\,q_{S}(y_{S^{c}}\mid x_{S^{c}};\,x_{S})}\right\}.$
Let $K_{S}$ be the resulting MH transition kernel:
$K_{S}(x,dy)\;=\;Q_{S}(x,dy)\,\alpha_{S}(x,y)\;+\;\Big(1-\int Q_{S}(x,dz)\,\alpha_{S}(x,z)\Big)\delta_{x}(dy).$
Then $K_{S}$ is reversible with respect to $\pi$, i.e.
$\pi(dx)\,K_{S}(x,dy)\;=\;\pi(dy)\,K_{S}(y,dx),$
and therefore $\pi$ is invariant for $K_{S}$.
Proof.
Write $x=(x_{S},x_{S^{c}})$ and $y=(y_{S},y_{S^{c}})$. Because the block $S$ is frozen, any proposed $y$ satisfies $y_{S}=x_{S}$. For such $x,y$,
$\pi(x)\,Q_{S}(x,dy)\,\alpha_{S}(x,y)=\pi(x)\,\delta_{x_{S}}(dy_{S})\,q_{S}(y_{S^{c}}\mid x_{S^{c}};x_{S})\,\min\!\left\{1,\frac{\pi(y)\,q_{S}(x_{S^{c}}\mid y_{S^{c}};x_{S})}{\pi(x)\,q_{S}(y_{S^{c}}\mid x_{S^{c}};x_{S})}\right\}.$
Using $\min\{a,b\}=\min\{b,a\}$ and noting that $x_{S}=y_{S}$ implies the same conditioning argument $x_{S}$ appears in both forward and reverse proposal densities, we obtain the standard MH symmetry:
$\pi(x)\,Q_{S}(x,dy)\,\alpha_{S}(x,y)=\pi(y)\,Q_{S}(y,dx)\,\alpha_{S}(y,x).$
Integrating both sides over measurable sets yields detailed balance for the “move” part. The “stay” part (the probability mass at $y=x$) matches on both sides by construction, completing detailed balance:
$\pi(dx)\,K_{S}(x,dy)=\pi(dy)\,K_{S}(y,dx).$
Hence $\pi$ is invariant for $K_{S}$. ∎
Corollary 6.1 (Parallel independent updates).
Fix a nonempty frozen subset $S\subset\{1,\ldots,N\}$ and write $S^{c}=\{1,\ldots,N\}\setminus S$. Suppose that, conditional on the frozen block $x_{S}$, the per–coordinate proposals for $i\in S^{c}$ factorize as
$q_{S}(y_{S^{c}}\mid x_{S^{c}};x_{S})\;=\;\prod_{i\in S^{c}}q_{S,i}(y_{i}\mid x_{i};x_{S}),$
and define the single-site MH kernels $K_{S,i}$ on $\mathbb{R}^{d}$ (conditional on $x_{S}$) by
$K_{S,i}(x_{i},dy_{i})=q_{S,i}(y_{i}\mid x_{i};x_{S})\,\alpha_{i}(x_{i},y_{i};x_{S})\,dy_{i}+\Big(1-\int q_{S,i}(z\mid x_{i};x_{S})\,\alpha_{i}(x_{i},z;x_{S})\,dz\Big)\delta_{x_{i}}(dy_{i}),$
with acceptance probability
$\alpha_{i}(x_{i},y_{i};x_{S})\;=\;\min\Big\{1,\,\frac{\rho(y_{i})\,q_{S,i}(x_{i}\mid y_{i};x_{S})}{\rho(x_{i})\,q_{S,i}(y_{i}\mid x_{i};x_{S})}\Big\}.$
Then, with $\pi=\rho^{\otimes N}$, the joint kernel
$\widetilde{K}_{S}(x,dy)\;=\;\delta_{x_{S}}(dy_{S})\,\prod_{i\in S^{c}}K_{S,i}(x_{i},dy_{i})$
is reversible with respect to $\pi$ (hence $\pi$-invariant).
Proof.
For each $i\in S^{c}$ and fixed $x_{S}$, the single-site MH kernel $K_{S,i}$ is reversible with respect to $\rho$:
$\rho(x_{i})\,K_{S,i}(x_{i},dy_{i})\;=\;\rho(y_{i})\,K_{S,i}(y_{i},dx_{i}).$
Using $\pi(dx)=\prod_{j=1}^{N}\rho(x_{j})\,dx_{j}$,
$\pi(dx)\,\widetilde{K}_{S}(x,dy)$ $=\Big(\prod_{j\in S}\rho(x_{j})\,dx_{j}\Big)\,\delta_{x_{S}}(dy_{S})\,\prod_{i\in S^{c}}\big[\rho(x_{i})\,dx_{i}\,K_{S,i}(x_{i},dy_{i})\big]$
$=\Big(\prod_{j\in S}\rho(y_{j})\,dy_{j}\Big)\,\delta_{y_{S}}(dx_{S})\,\prod_{i\in S^{c}}\big[\rho(y_{i})\,dy_{i}\,K_{S,i}(y_{i},dx_{i})\big]$
$=\pi(dy)\,\widetilde{K}_{S}(y,dx),$
using single-site detailed balance in the middle equality and the fact $x_{S}=y_{S}$ under the delta. Hence $\widetilde{K}_{S}$ is reversible w.r.t. $\pi$. ∎
A.7 Neal’s Funnel
In the most general form, the non-centered Neal’s Funnel is given by
[…]
for $i\in[d-1]$. The probability distribution can be given explicitly as
$\rho(x,y)\ =\ \left[\frac{1}{\sigma\sqrt{2\pi}}\exp\left(-\frac{x^{2}}{2\sigma^{2}}\right)\right]\left[\frac{1}{(\exp(x/2)\sqrt{2\pi})^{d-1}}\prod_{i=1}^{d-1}\exp\left(-\frac{y_{i}^{2}}{2\exp(x)}\right)\right].$ (114)
From this, it still follows that $\mathbb{E}[x]=\mathbb{E}[y_{i}]=0$ and $\mathbb{E}[x^{2}]=\sigma^{2}$.
To find $\mathbb{E}[y_{i}^{n}]$, note when $n=2k+1$ for $k\in\mathbb{N}$, $y_{i}^{2k+1}\rho(x,y)$ is an odd function in $y_{i}$, and thus $\mathbb{E}[y_{i}^{2k+1}]=0$.
When $n=2k$, the two integral formulas
$\int^{\infty}_{-\infty}e^{-ax^{2}-bx}dx\ =\ \sqrt{\frac{\pi}{a}}e^{b^{2}/4a},\qquad\text{and}\qquad\int^{\infty}_{0}x^{n}e^{-ax^{2}}\ =\ \frac{1}{2}a^{-(n+1)/2}\,\Gamma\left(\frac{n+1}{2}\right),$ (115)
for $a>0$, $n\in\mathbb{N}$, and $b\in\mathbb{E}$, yields, by direct computation
$\mathbb{E}[y_{i}^{2k}]\$$=\ \int_{\mathbb{R}^{d}}y_{i}^{2k}\rho(x,y)\,dx\,dy_{1}\,\ldots\,dy_{d-1}$ (116)
$=\ \frac{1}{2\pi\sigma}\int_{\mathbb{R}^{2}}y_{i}^{2k}\exp\left(-\frac{x^{2}}{2\sigma^{2}}-\frac{x}{2}\right)\exp\left(-\frac{y_{i}^{2}}{2\exp(x)}\right)\,dx\,dy_{i}.$ (117)
Since the integrated is even in $y_{i}$,
$\mathbb{E}[y_{i}^{2k}]\$$=\ \frac{1}{\pi\sigma}\int_{\mathbb{R}}\int_{0}^{\infty}y_{i}^{2k}\exp\left(-\frac{x^{2}}{2\sigma^{2}}-\frac{x}{2}\right)\exp\left(-\frac{y_{i}^{2}}{2\exp(x)}\right)\,dy_{i}\,dx$ (118)
$=\frac{1}{2\pi\sigma}\int_{\mathbb{R}}(2\exp(x))^{(2k+1)/2}\Gamma\left(\frac{2k+1}{2}\right)\exp\left(-\frac{x^{2}}{2\sigma^{2}}-\frac{x}{2}\right)\,dx$ (119)
$=\ \frac{2^{(2k+1)/2}}{2\pi\sigma}\Gamma\left(\frac{2k+1}{2}\right)\int_{\mathbb{R}}\exp\left(-\frac{x^{2}}{2\sigma^{2}}-\frac{x}{2}+\frac{(2k+1)x}{2}\right)\,dx$ (120)
$=\ \frac{2^{k+1/2}}{2\pi\sigma}\Gamma\left(\frac{2k+1}{2}\right)\int_{\mathbb{R}}\exp\left(-\frac{x^{2}}{2\sigma^{2}}+kx\right)\,dx$ (121)
$=\ \frac{2^{k+1/2}}{2\pi\sigma}\Gamma\left(\frac{2k+1}{2}\right)\sqrt{2\pi\sigma^{2}}e^{k^{2}\sigma^{2}/2}$ (122)
$=\ \frac{2^{k}}{\sqrt{\pi}}\Gamma\left(k+\frac{1}{2}\right)e^{k^{2}\sigma^{2}/2}.$ (123)
Using the known relationship
$\Gamma\left(k+\frac{1}{2}\right)\ =\ \frac{(2k)!}{4^{k}k!}\sqrt{\pi},$ (124)
we can thus simplify
$\mathbb{E}[y_{i}^{2k}]\ =\ \frac{2^{k}(2k)!}{4^{k}k!}e^{k^{2}\sigma^{2}/2}.$ (125)
Thus
$\mathbb{E}[y_{i}^{2}]\ =\ e^{\sigma^{2}/2},\qquad\text{and}\qquad\mathbb{E}[y_{i}^{4}]\ =\ 3e^{2\sigma^{2}}.$ (126)
Appendix B: Further details on posteriordb Experiment
This section of the Appendix illustrates the details of the posteriors used in Section 7.2. The details in Table 7 include the corresponding number of the posterior from the figures, the name of the model and datasets, and the number of dimensions of the sampling space. Note that the number of dimensions here corresponds to the number of unconstrained parameters as defined in the stan [8] model provided from posteriordb [28].
Table 7: Details of the posterior models from posteriordb
No. Name Dim No. Name Dim
$1.$ arK-arK 77 $24.$ kidiq_with_mom_work-kidscore_interaction_c2 55
$2.$ arma-arma11 44 $25.$ kidiq_with_mom_work-kidscore_interaction_z 55
$3.$ bball_drive_event_0-hmm_drive_0 66 $26.$ kidiq_with_mom_work-kidscore_mom_work 55
$4.$ bball_drive_event_1-hmm_drive_1 66 $27.$ kilpisjarvi_mod-kilpisjarvi 88
$5.$ diamonds-diamonds $26$ $28.$ low_dim_gauss_mix-low_dim_gauss_mix 55
$6.$ earnings-earn_height 33 $29.$ mesquite-logmesquite 88
$7.$ earnings-log10earn_height 33 $30.$ mesquite-logmesquite_logva 55
$8.$ earnings-logearn_height 33 $31.$ mesquite-logmesquite_logvas 88
$9.$ earnings-logearn_height_male 44 $32.$ mesquite-logmesquite_logvash 77
$10.$ earnings-logearn_interaction 55 $33.$ mesquite-logmesquite_logvolume 33
$11.$ earnings-logearn_interaction_z 55 $34.$ mesquite-mesquite 88
$12.$ earnings-logearn_logheight_male 44 $35.$ nes1972-nes $10$
$13.$ eight_schools-eight_schools_noncentered $10$ $36.$ nes1976-nes $10$
$14.$ garch-garch11 44 $37.$ nes1980-nes $10$
$15.$ gp_pois_regr-gp_pois_regr $13$ $38.$ nes1984-nes $10$
$16.$ gp_pois_regr-gp_regr 33 $39.$ nes1988-nes $10$
$17.$ hmm_example-hmm_example 44 $40.$ nes1992-nes $10$
$18.$ hudson_lynx_hare-lotka_volterra 88 $41.$ nes1996-nes $10$
$19.$ kidiq-kidscore_interaction 55 $42.$ nes2000-nes $10$
$20.$ kidiq-kidscore_momhs 55 $43.$ one_comp_mm_elim_abs-one_comp_mm_elim_abs 44
$21.$ kidiq-kidscore_momiq 44 $44.$ sblrc-blr 66
$22.$ kidiq-kidscore_momhsiq 33 $45.$ sblri-blr 66
$23.$ kidiq_with_mom_work-kidscore_interaction_c 55